### v1

# On the backward stability of s-step GMRES

Abstract
Communication, i.e., data movement, is a critical bottleneck for the performance of classical Krylov subspace method solvers on modern computer architectures. Variants of these methods which avoid communication have been introduced, which, while equivalent in exact arithmetic, can be unstable in finite precision. In this work, we address the backward stability of $s$-step GMRES, also known as communication-avoiding GMRES. We present a framework for simplifying the analysis of $s$-step GMRES, which includes standard GMRES ($s=1$) as a special case, by isolating the effects of rounding errors in the QR factorization and the solution of the least squares problem. Using this framework, we analyze $s$-step GMRES with popular block orthogonalization methods: block modified Gram–Schmidt and reorthogonalized block classical Gram–Schmidt algorithms.
An example illustrates the resulting instability of $s$-step GMRES when paired with the classical $s$-step Arnoldi process and shows the limitations of popular strategies for resolving this instability. To address this issue, we propose a modified Arnoldi process that allows for much larger block size $s$ while maintaining satisfactory accuracy, as confirmed by our numerical experiments.
Keywords: backward stability, $s$-step GMRES, communication-avoiding, Arnoldi process
MSC codes: 65F10, 65F50, 65G50
1 Introduction
Given a nonsingular square matrix $A\in\mathbb{R}^{n\times n}$ and a right-hand side $b\in\mathbb{R}^{n}$, this work considers the iterative solution of the linear system
$Ax=b.$ (1)
A popular method for this problem is the generalized minimal residual algorithm (GMRES) introduced in [20], which chooses $x^{(i)}\in x^{(0)}+\mathcal{K}_{i}(A,r)$ to minimize $\bigl{\|}Ax^{(i)}-b\bigr{\rVert}_{2}$ in the $i$-th iteration with a Krylov subspace $\mathcal{K}_{i}(A,r)=\operatorname{span}\{r,Ar,\dotsc,A^{i-1}r\}$ and the initial residual $r=b-Ax^{(0)}$.
The standard GMRES algorithm creates and orthogonalizes the Krylov basis one vector at a time, predominantly making use of BLAS 2 operations. On modern computer architectures, the performance of classical iterative solvers is heavily dominated by communication, i.e., data movement and synchronization. This motivated the introduction of $s$-step (also called communication-avoiding) variants of the GMRES algorithm, which can reduce asymptotic latency costs by a factor of $s$ and can take advantage of BLAS 3 operations. In [14, Section 3.6], it is demonstrated that CA-GMRES provides a speedup ranging from 1.31×1.31$\times$1.31 × to 4.3×4.3$\times$4.3 × compared to the standard GMRES on an Intel platform with 8 cores. The authors in [22] show that CA-GMRES can achieve speedups of up to 2×2$\times$2 × over the standard GMRES on one or multiple GPUs. Particularly, the orthogonalization phase in CA-GMRES can reach up to 4.16×4.16$\times$4.16 × speedup compared to the same phase in the standard GMRES when $s=10$ is used.
In each step, the $s$-step GMRES algorithm generates a block Krylov basis consisting of $s$ vectors formed as follows:
$\begin{bmatrix}p_{0}(A)v&p_{1}(A)v&\dotsc&p_{s-1}(A)v\end{bmatrix},$ (2)
[…]
Considering rounding errors, the backward stability of the standard GMRES algorithms with Householder and modified Gram–Schmidt orthogonalization (MGS) has been investigated in [12, 19]. Recently, [5] introduced a comprehensive framework to simplify the rounding error analysis of GMRES algorithms. Building on this framework, we analyze the rounding error of the $s$-step GMRES algorithm to illustrate how the condition number of the basis affects the backward error in this study. Our analysis is performed in a way so that it is possible identify the sources of errors, and clarifying the impact of different orthogonalization methods on the backward error of the computed $s$-step GMRES results and broadening the applicability of the framework introduced by [5]. In addition, our analysis formally shows that the backward error of $s$-step GMRES is influenced by the condition number of the basis. From this analysis, we further propose a modified Arnoldi process to allow for the use of a substantially larger block size $s$, as confirmed by numerical experiments, increasing the cost of the orthogonalization by a factor of 2.
The reminder of this paper is organized as follows. In Section 2, we give an overview of the $s$-step Arnoldi and GMRES algorithms. In Section 3, we present our abstract framework for the analysis of the $s$-step GMRES method. In Section 4, we invoke our abstract framework to analyze the backward error of the $s$-step GMRES algorithm with different commonly-used block orthogonalization methods. We discuss the implications of the theoretical results and comment on stopping criteria in Section 5. Then we propose a modified $s$-step Arnoldi process in Section 6 to improve the backward stability of the $s$-step GMRES. Numerical experiments are presented in Section 7 which compare the $s$-step GMRES algorithm with the modified and classical $s$-step Arnoldi processes.
[…]
2 $s$-step Arnoldi and GMRES algorithms
In this section, we begin with an overview of the $s$-step Arnoldi and GMRES algorithms without employing any preconditioners, followed by a discussion of their variants incorporating left and right preconditioners.
The standard GMRES algorithm, namely $s$-step GMRES algorithm with $s=1$, employs the Arnoldi process to build an orthonormal basis $V_{1:i+1}$ of the Krylov subspace $\mathcal{K}_{i+1}(A,r)$ with $r=b-Ax^{(0)}$, satisfying
$\begin{bmatrix}r&W_{1:i}\end{bmatrix}=V_{1:i+1}R_{1:i+1}$
with $W_{1:i}=AV_{1:i}$ and the upper triangular matrix $R_{1:i+1}$, after $i$ iterations, and then update $x^{(i)}\leftarrow x^{(0)}+V_{1:i}y^{(i)}$ if converged, where $y^{(i)}=\operatorname*{\arg\min}_{y}\bigl{\|}\beta e_{1}-H_{1:i+1,1:i}y\bigr{\rVert}$ with $\beta=\bigl{\|}r\bigr{\rVert}$, $e_{1}$ denoting the first column of the $(i+1)$-by-$(i+1)$ identity matrix, and $H_{1:i+1,1:i}=R_{1:i+1,2:i+1}$.
When $s>1$, the $s$-step GMRES algorithm handles $s$ vectors at a time in every iteration. Specifically, in the $i$-th iteration, the $s$-step algorithm first builds an orthonormal basis $V_{1:is+1}$, by the $s$-step Arnoldi process, satisfying
$\begin{bmatrix}r&W_{1:is}\end{bmatrix}=V_{1:is+1}R_{1:is+1},$
where $W_{1:is}=AB_{1:is}$ and each subblock of $B_{1:is}$, namely $B_{(k-1)s+1:ks}$ with $k=1,2,\ldots,i$, satisfies
$B_{(k-1)s+1:ks}=\begin{bmatrix}p_{0}(A)V_{(k-1)s+1}&p_{1}(A)V_{(k-1)s+1}&\dotsb&p_{s-1}(A)V_{(k-1)s+1}\end{bmatrix}$
with given polynomials $p_{0}$, $p_{1}$, $\ldots$, $p_{s-1}$. Next, the $s$-step GMRES algorithm updates $x^{(i)}$ by $x^{(0)}+B_{1:is}y^{(i)}$, where $y^{(i)}=\operatorname*{\arg\min}_{y}\bigl{\|}\beta e_{1}-H_{1:is+1,1:is}y\bigr{\rVert}$ with $H_{1:is+1,1:is}=R_{1:is+1,2:is+1}$. Notice that only the subdiagonal elements are nonzero in the lower triangular part of $H_{1:is+1,1:is}$. The least squares problem $\min_{y}\bigl{\|}\beta e_{1}-H_{1:is+1,1:is}y\bigr{\rVert}$ is often solved by applying the Givens QR factorization to $H_{1:is+1,1:is}$ such that
$H_{1:is+1,1:is}=G_{1:is+1}T_{1:is+1,1:is}$
with an orthogonal matrix $G_{1:is+1}$ and upper triangular matrix $T_{1:is}$. Then $y^{(i)}$ can be computed by solving the triangular system $T_{1:is}y^{(i)}=\beta G_{1,1:is}^{\top}$, since
$y^{(i)}$ $=\operatorname*{\arg\min}_{y}\bigl{\|}\beta e_{1}-G_{1:is+1}T_{1:is+1,1:is}y\bigr{\rVert}$
$=\operatorname*{\arg\min}_{y}\bigl{\|}\beta G_{1,1:is+1}^{\top}-T_{1:is+1,1:is}y\bigr{\rVert}$
$=\operatorname*{\arg\min}_{y}\bigl{\|}\beta G_{1,1:is}^{\top}-T_{1:is}y\bigr{\rVert}.$
This $s$-step GMRES algorithm is the so-called “non-traditional” variant introduced by [15].
Furthermore, if we consider the left and right preconditioners, i.e., $M_{L}$ and $M_{R}$, $s$-step Arnoldi process aims to build an orthonormal basis $V_{1:is+1}$ of $\mathcal{K}_{i+1}(M_{L}^{-1}AM_{R}^{-1},r)$ with $r=M_{L}^{-1}(b-Ax^{(0)})$. This means that $V_{1:is+1}$ satisfies
$\begin{bmatrix}r&W_{1:is}\end{bmatrix}=V_{1:is+1}R_{1:is+1}$
with $W_{1:is}=M_{L}^{-1}AM_{R}^{-1}B_{1:is}$. Then $s$-step GMRES updates $x^{(i)}$ through $x^{(0)}+M_{R}^{-1}B_{1:is}y^{(i)}$, where $y^{(i)}$ is the optimal solution for $\min_{y}\bigl{\|}\beta e_{1}-H_{1:is+1,1:is}y\bigr{\rVert}$ with $\beta=\bigl{\|}r\bigr{\rVert}$ and $H_{1:is+1,1:is}=R_{1:is+1,2:is+1}$.
We summarize the above preconditioned $s$-step Arnoldi and GMRES algorithms in Algorithms 1 and 2, respectively.
0:  A matrix $A\in\mathbb{R}^{n\times n}$, a vector $r$, a block size $s$, a left-preconditioner $M_{L}\in\mathbb{R}^{n\times n}$, a right-preconditioner $M_{R}\in\mathbb{R}^{n\times n}$, the basis $B_{1:(i-1)s}$ and the preconditioned basis $Z_{1:(i-1)s}$ generated by the first $i-1$ classical Arnoldi steps, the matrix $W_{1:(i-1)s}$, the orthonormal matrix $V_{1:(i-1)s+1}$, and the upper triangular matrix $R_{1:(i-1)s+1}$ satisfying $\begin{bmatrix}r&W_{1:(i-1)s}\end{bmatrix}=V_{1:(i-1)s+1}R_{1:(i-1)s+1}$.
0:  The basis $B_{1:is}$, the preconditioned basis $Z_{1:is}$, the matrices $W_{1:is}$, $V_{1:is+1}$, and $R_{1:is+1}$ satisfying $\begin{bmatrix}r&W_{1:is}\end{bmatrix}=V_{1:is+1}R_{1:is+1}$.
1:  $B_{(i-1)s+1:is}\leftarrow\begin{bmatrix}p_{0}(A)V_{(i-1)s+1}&p_{1}(A)V_{(i-1)s+1}&\dotsi&p_{s-1}(A)V_{(i-1)s+1}\end{bmatrix}$.
2:  $Z_{(i-1)s+1:is}\leftarrow M_{R}^{-1}B_{(i-1)s+1:is}$
3:  $W_{(i-1)s+1:is}\leftarrow M_{L}^{-1}AZ_{(i-1)s+1:is}$.
4:  Compute the QR factorization of $\begin{bmatrix}r&W_{1:is}\end{bmatrix}=V_{1:is+1}R_{1:is+1}$ based on $\begin{bmatrix}r&W_{1:(i-1)s}\end{bmatrix}=V_{1:(i-1)s+1}R_{1:(i-1)s+1}$.
Algorithm 1 The $i$-th iteration of the $s$-step Arnoldi process
0:  A matrix $A\in\mathbb{R}^{n\times n}$, a right-hand side $b\in\mathbb{R}^{n}$, an initial guess $x^{(0)}\in\mathbb{R}^{n}$, a block size $s$, a left-preconditioner $M_{L}\in\mathbb{R}^{n\times n}$, and a right-preconditioner $M_{R}\in\mathbb{R}^{n\times n}$.
0:  A computed solution $x\in\mathbb{R}^{n}$ approximating the solution of $Ax=b$.
1:  $r\leftarrow M_{L}^{-1}(b-Ax^{(0)})$ and $\beta\leftarrow\beta=\bigl{\|}r\bigr{\rVert}$.
2:  $V_{1}=r/\beta$.
3:  for $i=1:n/s$ do
4:     Perform the $i$-th step of the $s$-step Arnoldi process (e.g., Algorithm 1) to obtain the basis $B_{1:is}$, the preconditioned basis $Z_{1:is}$, the orthonormal matrix $V_{1:is+1}$, and the upper triangular matrix $R_{1:is+1}$ satisfying $\begin{bmatrix}r&W_{1:is}\end{bmatrix}=V_{1:is+1}R_{1:is+1}$ with $W_{1:is}\leftarrow M_{L}^{-1}AZ_{1:is}$.
5:     $H_{1:(i-1)s+1,1:(i-1)s}\leftarrow R_{1:(i-1)s+1,2:(i-1)s+1}$.
6:     Compute the QR factorization $H_{1:is+1,(i-1)s+1:is}=G_{1:is+1}T_{1:is+1,(i-1)s+1:is}$ by Givens rotations with an orthogonal matrix $G_{1:is+1}$, based on $H_{1:(i-1)s+1,1:(i-1)s}=G_{1:(i-1)s+1}T_{1:(i-1)s+1,1:(i-1)s}$.
7:     if the stopping criterion is satisfied then
8:        Solve the triangular system $T_{1:is}y^{(i)}=\beta G_{1,1:is}^{\top}$ to obtain $y^{(i)}\in\mathbb{R}^{is}$.
9:        return  $x=x^{(i)}\leftarrow x^{(0)}+Z_{1:is}y^{(i)}$.
10:     end if
11:  end for
Algorithm 2 The $s$-step GMRES algorithm
3 An abstract framework for backward stability of $s$-step GMRES
In the recent work [5], the authors propose a framework to study the backward stability of what they call “modular GMRES” as follows:
  1. 1.
    Compute $Z_{1:(i+1)s}=M_{R}^{-1}V_{1:(i+1)s}$ and $W_{1:(i+1)s}=M_{L}^{-1}AZ_{1:(i+1)s}$.
  2. 2.
    Solve $y^{(i+1)}=\operatorname*{\arg\min}_{y}\bigl{\|}M_{L}^{-1}b-W_{1:(i+1)s}y\bigr{\rVert}$.
  3. 3.
    Compute the solution approximation $x^{(i+1)}=Z_{1:(i+1)s}y^{(i+1)}$.
To assess the backward stability of GMRES with various orthogonalization methods, it is necessary to evaluate the rounding errors in Step 2 for the chosen orthogonalization method within this framework. Note that Step 2 in modular GMRES comprises orthogonalization, Givens QR factorization, and solving a triangular system. Therefore, an analysis of all three parts is required when modifying the orthogonalization method used in GMRES, which is less convenient and unnecessary, and further causes ambiguity regarding the requirements of the orthogonalization method.
In this section, we extend the above framework to $Z_{1:(i+1)s}=M_{R}^{-1}B_{1:(i+1)s}$ focusing on the backward stability of the $s$-step GMRES method. This includes Algorithm 2 with the classical $s$-step Arnoldi process (Algorithm 1), and the standard GMRES algorithm, which is essentially a specific case of $s$-step GMRES with $s=1$ and $p_{0}(A)=I$, hence $B_{1:(i+1)s}=V_{1:(i+1)s}$. Furthermore, we delve into the block orthogonalization, Givens QR decomposition, and triangular system solving steps, which are actually performed in the algorithm, as opposed to analyzing $y^{(i+1)}=\operatorname*{\arg\min}_{y}\bigl{\|}M_{L}^{-1}b-W_{1:(i+1)s}y\bigr{\rVert}$. From this analysis, we will present the assumptions only relevant to the block orthogonalization process itself. Consequently, we only need to consider the rounding errors in the orthogonalization process when modifying the orthogonalization method utilized in the GMRES algorithm. We note that the approach here thus represents an improved version of the modular GMRES framework in [5] for the standard (non $s$-step) GMRES case.
According to Algorithm 2, the first $(i+1)$ iterations can be summarized as the following four steps. Without loss of generality, we assume $x^{(0)}=0$ for simplicity in the following analysis.
  1. 1.
    Compute $Z_{1:(i+1)s}=M_{R}^{-1}B_{1:(i+1)s}$ and $W_{1:(i+1)s}=M_{L}^{-1}AZ_{1:(i+1)s}$.
  2. 2.
    Compute the QR factorization
    $\begin{bmatrix}r&W_{1:(i+1)s}\end{bmatrix}=V_{1:(i+1)s+1}R_{1:(i+1)s+1},$
    where $V_{1:(i+1)s+1}$ is orthonormal, and $R_{1:(i+1)s+1}$ is upper triangular.
  3. 3.
    Solve the least squares problem $y^{(i+1)}=\operatorname*{\arg\min}_{y}\bigg{\|}R_{1:(i+1)s+1}\begin{bmatrix}1\\ -y\end{bmatrix}\bigg{\rVert}$, which is through solving the triangular system $T_{1:(i+1)s}y^{(i+1)}=\beta G_{1,1:(i+1)s}^{\top}$ to obtain $y^{(i+1)}$, where $T_{1:(i+1)s+1,1:(i+1)s}$ is $R$-factor of the QR factorization by Givens rotations, i.e.,
    $H_{1:(i+1)s+1,1:(i+1)s}=G_{1:(i+1)s+1}T_{1:(i+1)s+1,1:(i+1)s}$
    with $H_{1:(i+1)s+1,1:(i+1)s}=R_{1:(i+1)s+1,2:(i+1)s+1}$ and an orthogonal matrix $G_{1:(i+1)s+1}$.
  4. 4.
    Compute the solution approximation $x^{(i+1)}=Z_{1:(i+1)s}y^{(i+1)}$.
Taking rounding errors into account, we further assume that each line satisfies the following, to be described below. In the remainder of this subsection, $\varepsilon_{*}$, $\delta_{*}$, and $\omega_{*}$ denote constants in the rounding error analysis of different operations, which helps to identify the source of errors in the final results. We use $\delta$ to indicate constants associated with the left and right preconditioners, and use $\omega$ to denote constants relevant to the block orthogonalization process. Constants that are independent of these two parts are denoted by $\varepsilon$, which can be considered as $\varepsilon_{*}\in(0,1)$, as ensured by standard rounding error analysis; see, e.g., [13].
Step 1: Generating the basis
This step satisfies
$\widehat{Z}_{1:(i+1)s}$ $=M_{R}^{-1}\hat{B}_{1:(i+1)s}+\Delta Z_{1:(i+1)s},\quad\bigl{\|}\Delta Z_{j}\bigr{\rVert}\leq\delta_{M_{R}^{-1}}\bigl{\|}M_{R}^{-1}\bigr{\rVert}\bigl{\|}\hat{B}_{j}\bigr{\rVert},$ (3)
$\widehat{W}_{1:(i+1)s}$ $=M_{L}^{-1}(A\widehat{Z}_{1:(i+1)s}+\Delta C_{1:(i+1)s})+\Delta D_{1:(i+1)s}.$ (4)
Here $\Delta C_{1:(i+1)s}$ and $\Delta D_{1:(i+1)s}$ come from, respectively, computing $A\widehat{Z}_{1:(i+1)s}$ and applying $M_{L}^{-1}$ to $A\widehat{Z}_{1:(i+1)s}+\Delta C_{1:(i+1)s}$, which satisfy
$\bigl{\|}\Delta C_{j}\bigr{\rVert}$ $\leq\varepsilon_{AZ}\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\bigl{\|}\widehat{Z}_{j}\bigr{\rVert},\quad\bigl{\|}\Delta D_{j}\bigr{\rVert}\leq\delta_{M_{L}^{-1}AZ}\bigl{\|}M_{L}^{-1}\bigr{\rVert}\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\bigl{\|}\widehat{Z}_{j}\bigr{\rVert}$
for any $j\in\{1,2,\dotsc,(i+1)s\}$. Let $\Delta W_{1:(i+1)s}=M_{L}^{-1}\Delta C_{1:(i+1)s}+\Delta D_{1:(i+1)s}$. $\widehat{W}_{1:(i+1)s}$ further satisfies
$\widehat{W}_{1:(i+1)s}=M_{L}^{-1}A\widehat{Z}_{1:(i+1)s}+\Delta W_{1:(i+1)s},$ (5)
with $\bigl{\|}\Delta W_{j}\bigr{\rVert}\leq(\varepsilon_{AZ}+\delta_{M_{L}^{-1}AZ})\bigl{\|}M_{L}^{-1}\bigr{\rVert}\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\bigl{\|}\widehat{Z}_{j}\bigr{\rVert}$, which means
$\bigl{\|}\Delta W_{1:(i+1)s}D_{1:(i+1)s}^{-1}\bigr{\rVert}_{\mathsf{F}}\leq(\varepsilon_{AZ}+\delta_{M_{L}^{-1}AZ})\bigl{\|}M_{L}^{-1}\bigr{\rVert}\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\bigl{\|}\tilde{Z}_{1:(i+1)s}\bigr{\rVert}_{\mathsf{F}},$ (6)
where $\widehat{Z}_{1:(i+1)s}=\tilde{Z}_{1:(i+1)s}D_{1:(i+1)s}$ with a diagonal positive definite matrix $D_{1:(i+1)s}$.
Step 2: Backward stability of the block orthogonalization
Let $r=M_{L}^{-1}(b-Ax^{(0)})=M_{L}^{-1}b$ with $x^{(0)}=0$ and $\beta=\bigl{\|}\widehat{r}\bigr{\rVert}$ satisfy
$\widehat{r}=M_{L}^{-1}b+\Delta r,\quad\bigl{\|}\Delta r\bigr{\rVert}\leq\delta_{M_{L}^{-1}b}\bigl{\|}M_{L}^{-1}\bigr{\rVert}\bigl{\|}b\bigr{\rVert},$ (7)
$\widehat{\beta}=(1+\varepsilon_{r})\beta.$ (8)
Then we assume the backward stability, which is satisfied by most commonly-used block orthogonalization schemes summarized in [10], of the block orthogonalization method used in Step 2; i.e., for any $\phi>0$,
$\begin{bmatrix}\phi\widehat{r}&\widehat{W}_{1:(i+1)s}\end{bmatrix}+\Delta E_{1:(i+1)s+1}(\phi)=\bar{V}_{1:(i+1)s+1}\widehat{R}_{1:(i+1)s+1}(\phi),$ (9)
where $\widehat{R}_{1:(i+1)s+1}(\phi)=\begin{bmatrix}\widehat{R}_{1:(i+1)s+1,1}(\phi)&\widehat{R}_{1:(i+1)s+1,2:(i+1)s+1}\end{bmatrix}$ and the backward error $\Delta E_{1:(i+1)s+1}(\phi)=\begin{bmatrix}\Delta E_{1}(\phi)&\Delta E_{2:(i+1)s+1}\end{bmatrix}$ satisfies
$\bigl{\|}\Delta E_{1}(\phi)\bigr{\rVert}\leq\omega_{\mathrm{qr}}\bigl{\|}\phi\widehat{r}\bigr{\rVert}=\omega_{\mathrm{qr}}\phi\beta\quad\text{and}\quad\bigl{\|}\Delta E_{j}\bigr{\rVert}\leq\omega_{\mathrm{qr}}\bigl{\|}\widehat{W}_{j-1}\bigr{\rVert}$ (10)
for any $j\in\{2,\dotsc,(i+1)s+1\}$, which implies
$\begin{split}&\begin{split}&\bigg{\|}\Delta E_{1:is+j}(\phi)\begin{bmatrix}1&0\\ 0&D_{1:is+j-1}^{-1}\end{bmatrix}\bigg{\rVert}_{\mathsf{F}}\\ &\quad\leq\omega_{\mathrm{qr}}\phi\beta+\omega_{\mathrm{qr}}\bigl{\|}\widehat{W}_{1:is+j}D_{1:is+j-1}^{-1}\bigr{\rVert}_{\mathsf{F}}\\ &\quad\leq\omega_{\mathrm{qr}}\phi\beta+\omega_{\mathrm{qr}}\bigl{(}1+\varepsilon_{AZ}+\delta_{M_{L}^{-1}AZ}\bigr{)}\bigl{\|}M_{L}^{-1}\bigr{\rVert}\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\bigl{\|}\tilde{Z}_{is+j-1}\bigr{\rVert}_{\mathsf{F}},\\ \end{split}\\ &\begin{split}&\bigg{\|}\Delta E_{1:(i+1)s+1}(\phi)\begin{bmatrix}1\\ -\widehat{y}^{(i+1)}\end{bmatrix}\bigg{\rVert}\\ &\quad\leq\bigg{\|}\Delta E_{1:(i+1)s+1}(\phi)\begin{bmatrix}1&0\\ 0&D_{1:(i+1)s}^{-1}\end{bmatrix}\begin{bmatrix}1\\ -D_{1:(i+1)s}\widehat{y}^{(i+1)}\end{bmatrix}\bigg{\rVert}\\ &\quad\leq\omega_{\mathrm{qr}}\phi\beta+\omega_{\mathrm{qr}}\bigl{(}1+\varepsilon_{AZ}+\delta_{M_{L}^{-1}AZ}\bigr{)}\bigl{\|}M_{L}^{-1}\bigr{\rVert}\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\bigl{\|}\tilde{Z}_{1:(i+1)s}\bigr{\rVert}_{\mathsf{F}}\bigl{\|}D_{1:(i+1)s}\widehat{y}^{(i+1)}\bigr{\rVert},\end{split}\end{split}$ (11)
for any $j\in\{1,2,\dotsc,s\}$. Here $\bar{V}_{*}$ can be the computed result $\widehat{V}_{*}$ if using the reorthogonalized block classical Gram–Schmidt algorithm, or an exact orthonormal matrix $\tilde{V}_{*}$ if using the block modified Gram–Schmidt algorithm.
In addition, we define $\omega_{is+j}$ to describe the loss of orthogonality of the $Q$-factor, i.e., for any $j\in\{1,2,\dotsc,s\}$,
$\begin{split}&\bigl{\|}\bar{V}_{1:is+j}^{\top}\bar{V}_{1:is+j}-I\bigr{\rVert}_{\mathsf{F}}\leq\omega_{is+j},\\ &\bigl{\|}\bar{V}_{1:is+j}\bigr{\rVert}\leq\sqrt{\bigl{\|}I\bigr{\rVert}+\bigl{\|}\bar{V}_{1:is+j}^{\top}\bar{V}_{1:is+j}-I\bigr{\rVert}}\leq\sqrt{1+\omega_{is+j}}.\end{split}$ (12)
Step 2: Loss of orthogonality in the block orthogonalization
In this part, we discuss the iteration where $\widehat{V}_{*}$ is no longer well-conditioned, which is equivalent to the condition that the columns of $\widehat{V}_{*}$ have lost orthogonality. This iteration is called the “key dimension” in [5], since the loss of orthogonality always happens when the condition number of $\begin{bmatrix}\phi r&\widehat{W}_{1:p}D_{1:p}^{-1}\end{bmatrix}$ becomes large enough. We employ a similar definition of the key dimension introduced by [5] here. We define $p=i^{*}s+j^{*}$ to be the iteration in which we reach the key dimension, i.e., when we have
$\sigma_{\min}\Big{(}\begin{bmatrix}\phi\widehat{r}&\widehat{W}_{1:p}D_{1:p}^{-1}\end{bmatrix}\Big{)}\leq\omega_{\kappa(r,W)}\Big{\|}\begin{bmatrix}\phi\widehat{r}&\widehat{W}_{1:p}D_{1:p}^{-1}\end{bmatrix}\Big{\rVert}_{\mathsf{F}}.$ (13)
Together with
$\sigma_{\min}(\widehat{W}_{1:p}D_{1:p}^{-1})\geq\omega_{\kappa(W)}\bigl{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\bigr{\rVert}_{\mathsf{F}},$ (14)
(13) indicates that $\widehat{r}$ lies in the range of $\widehat{W}_{1:p}$.
Step 3: Solving the least squares problem
We only need to consider the case of the key dimension. The procedure of solving the least squares problem in Step 3 at the key dimension satisfies
$\widehat{y}^{(k)}$ $=\operatorname*{\arg\min}_{y}\bigg{\|}(\widehat{R}_{1:p+1}+\Delta R_{1:p+1})\begin{bmatrix}1\\ -y\end{bmatrix}\bigg{\rVert},\quad\bigl{\|}\Delta R_{1:p+1,j}\bigr{\rVert}\leq\varepsilon_{ls}\bigl{\|}\widehat{R}_{1:p+1,j}\bigr{\rVert},$ (15)
for any $j=1,\dotsc,p+1$. The error $\Delta R_{1:p+1}$ further satisfies
$\Big{\|}\begin{bmatrix}\phi\Delta R_{1:p+1,1}&\Delta R_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\Big{\rVert}_{\mathsf{F}}\leq\varepsilon_{ls}\Big{\|}\begin{bmatrix}\phi\widehat{R}_{1:p+1,1}&\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\Big{\rVert}_{\mathsf{F}}.$ (16)
About the least squares problem, [5] measures $\min_{y}{\bigl{\|}\widehat{r}-\widehat{W}_{1:p}y\bigr{\rVert}}$. It is essential to note that we avoid accessing $\min_{y}{\bigl{\|}\widehat{r}-\widehat{W}_{1:p}y\bigr{\rVert}}$, due to the substantial impact of the block orthogonalization process on this residual. In this context, we aim to distinguish the errors introduced by Steps 2 and 3 separately. Therefore, our attention is on (15), which is performed in Step 3 and remains unaffected by the block orthogonalization.
Step 4: Updating $x$
The step of updating $x$ satisfies
$\widehat{x}^{(i^{*})}=\widehat{Z}_{1:p}\widehat{y}^{(i^{*})}+\Delta x^{(i^{*})},\quad\bigl{\|}\Delta x^{(i^{*})}\bigr{\rVert}\leq\varepsilon_{mv}\bigl{\|}\tilde{Z}_{1:p}\bigr{\rVert}_{\mathsf{F}}\bigl{\|}D_{1:p}\widehat{y}^{(i^{*})}\bigr{\rVert}.$ (17)
Remark 1.
From standard rounding error analysis and [13, Equation (3.12)], we have $\varepsilon_{r}$, $\varepsilon_{AZ}$, $\varepsilon_{mv}=O(\bm{u})$. In [5, Equation (A.7)], the authors have given $\varepsilon_{ls}=O(\bm{u})$, which is from Givens QR factorization and the triangular solver.
The terms $\delta_{M_{L}^{-1}AZ}$, $\delta_{M_{L}^{-1}b}$, and $\delta_{M_{R}^{-1}}$ are highly dependent on the preconditioners $M_{L}$ and $M_{R}$. The remaining terms $\omega_{\mathrm{qr}}$, $\omega_{is+j}$, $\omega_{\kappa(r,W)}$, and $\omega_{\kappa(W)}$ are determined by the block orthogonalization algorithm employed in Step 2.
In the implementation of the $s$-step GMRES algorithm, the stopping criterion typically involves checking whether $\min_{y}\bigg{\|}(\widehat{R}_{1:is+j+1}+\Delta R_{1:is+j+1})\begin{bmatrix}1\\ -y\end{bmatrix}\bigg{\rVert}$ is sufficiently small. Thus, we first show that the residual $\bigg{\|}\widehat{R}_{1:p+1}\begin{bmatrix}1\\ -\widehat{y}^{(k)}\end{bmatrix}\bigg{\rVert}$ is small enough in Lemma 3. Then we prove that the backward error $\bigl{\|}b-A\widehat{x}^{(k)}\bigr{\rVert}$ can be bounded mainly by the residual of the least squares problem in Theorem 1.
To bound $\bigg{\|}\widehat{R}_{1:p+1}\begin{bmatrix}1\\ -\widehat{y}^{(k)}\end{bmatrix}\bigg{\rVert}$, we first need to establish the relationship between the result and the assumptions (13), (14), which amounts to bounding
$\sigma_{\min}\bigg{(}\begin{bmatrix}\phi\widehat{R}_{1:p+1,1}&\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\bigg{)}\quad\text{and}\quad\Big{\|}\begin{bmatrix}\phi\widehat{R}_{1:p+1,1}&\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\Big{\rVert}_{\mathsf{F}}$
by $\Big{\|}\begin{bmatrix}\phi r,\widehat{W}_{1:p}D_{1:p}^{-1}\end{bmatrix}\Big{\rVert}_{\mathsf{F}}$ and $\bigl{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\bigr{\rVert}_{\mathsf{F}}$. We summarize the result in the following lemma.
Lemma 1.
Assume that $\begin{bmatrix}\phi r,\widehat{W}_{1:p}D_{1:p}^{-1}\end{bmatrix}$ and $\widehat{W}_{1:p}D_{1:p}^{-1}$ satisfy (13) and (14), respectively. If also assuming $\bigl{\|}\bar{V}_{1:p}^{\top}\bar{V}_{1:p}-I\bigr{\rVert}_{\mathsf{F}}\leq\omega_{p}\leq 1/2$, then
$\sigma_{\min}\Big{(}\begin{bmatrix}\phi\widehat{R}_{1:p+1,1},\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\Big{)}\leq{}$ $\alpha_{11}\Big{\|}\begin{bmatrix}\phi r,\widehat{W}_{1:p}D_{1:p}^{-1}\end{bmatrix}\Big{\rVert}_{\mathsf{F}}+\alpha_{12}\bigl{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\bigr{\rVert}_{\mathsf{F}},$ (18)
$\Big{\|}\begin{bmatrix}\phi\widehat{R}_{1:p+1,1},\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\Big{\rVert}_{\mathsf{F}}\leq{}$ $\alpha_{21}\Big{\|}\begin{bmatrix}\phi r,\widehat{W}_{1:p}D_{1:p}^{-1}\end{bmatrix}\Big{\rVert}_{\mathsf{F}}+\alpha_{22}\bigl{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\bigr{\rVert}_{\mathsf{F}},$ (19)
$\sigma_{\min}(\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1})\geq{}$ $\alpha_{3}\bigl{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\bigr{\rVert}_{\mathsf{F}},$ (20)
$\bigl{\|}\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\bigr{\rVert}_{\mathsf{F}}\leq{}$ $\alpha_{4}\bigl{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\bigr{\rVert}_{\mathsf{F}},$ (21)
where
$\alpha_{11}=3(\omega_{\kappa(r,W)}+\omega_{\mathrm{qr}}),\quad\alpha_{12}=4\cdot\omega_{H},$ (22)
$\alpha_{21}=3(1+\omega_{\mathrm{qr}}),\quad\alpha_{22}=4\cdot\omega_{H},$ (23)
$\alpha_{3}=\frac{2}{3}(\omega_{\kappa(W)}-\omega_{\mathrm{qr}}-\omega_{H})-\omega_{H},$ (24)
$\alpha_{4}=3(1+\omega_{\mathrm{qr}}+\omega_{H})+\omega_{H},$ (25)
with $\omega_{H}$ defined by
$\max\bigl{(}\bigl{\lvert}\widehat{R}_{p+1,p+1}D_{p,p}^{-1}\bigr{\rvert},\bigl{\|}\bar{V}_{p+1}\widehat{R}_{p+1,2:p+1}D_{1:p}^{-1}\bigr{\rVert}_{\mathsf{F}}\bigr{)}\leq\omega_{H}\bigl{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\bigr{\rVert}_{\mathsf{F}},$ (26)
if $\bigl{\|}\bar{V}_{1:p}^{\top}\bar{V}_{1:p}-I\bigr{\rVert}_{\mathsf{F}}>0$; otherwise, $\omega_{H}=0$.
Before proving Lemma 1, we give the following lemma which will be used in the following proofs.
Lemma 2.
Given $X\in\mathbb{R}^{n\times m}$. Assume that $X+\Delta X=QU$, where $Q\in\mathbb{R}^{n\times m}$ satisfies $\bigl{\|}Q^{\top}Q-I\bigr{\rVert}_{\mathsf{F}}<1$. Then
$\bigl{\|}U\bigr{\rVert}_{\mathsf{F}}\leq\frac{\bigl{\|}Q\bigr{\rVert}(\bigl{\|}X\bigr{\rVert}_{\mathsf{F}}+\bigl{\|}\Delta X\bigr{\rVert}_{\mathsf{F}})}{1-\bigl{\|}Q^{\top}Q-I\bigr{\rVert}_{\mathsf{F}}}\quad\text{and}\quad\sigma_{\min}(U)\geq\frac{\sigma_{\min}(X)-\bigl{\|}\Delta X\bigr{\rVert}_{\mathsf{F}}}{\bigl{\|}Q\bigr{\rVert}}.$ (27)
Proof.
By $X+\Delta X=QU$, we have $U=Q^{\top}X+Q^{\top}\Delta X-(Q^{\top}Q-I)U$ and then
$\bigl{\|}U\bigr{\rVert}_{\mathsf{F}}\leq\bigl{\|}Q\bigr{\rVert}(\bigl{\|}X\bigr{\rVert}_{\mathsf{F}}+\bigl{\|}\Delta X\bigr{\rVert}_{\mathsf{F}})+\bigl{\|}Q^{\top}Q-I\bigr{\rVert}\bigl{\|}U\bigr{\rVert}_{\mathsf{F}},$ (28)
which gives the bound of $\bigl{\|}U\bigr{\rVert}_{\mathsf{F}}$. The bound of $\sigma_{\min}(U)$ has been proved in [7, Lemma 2]. ∎
We now prove Lemma 1.
Proof of Lemma 1.
From (9), if $\bar{V}_{1:p+1}$ is exact orthonormal, we have
$\sigma_{\min}\Big{(}\begin{bmatrix}\phi\widehat{R}_{1:p+1,1}&\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\Big{)}$ $\leq\sigma_{\min}\Big{(}\begin{bmatrix}\phi r,\widehat{W}_{1:p}D_{1:p}^{-1}\end{bmatrix}\Big{)}$
$\quad+\omega_{\mathrm{qr}}\Big{\|}\begin{bmatrix}\phi r,\widehat{W}_{1:p}D_{1:p}^{-1}\end{bmatrix}\Big{\rVert}_{\mathsf{F}},$
$\Big{\|}\begin{bmatrix}\phi\widehat{R}_{1:p+1,1}&\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\Big{\rVert}_{\mathsf{F}}$ $\leq(1+\omega_{\mathrm{qr}})\Big{\|}\begin{bmatrix}\phi r,\widehat{W}_{1:p}D_{1:p}^{-1}\end{bmatrix}\Big{\rVert}_{\mathsf{F}},$
$\sigma_{\min}(\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1})$ $\geq\sigma_{\min}(\widehat{W}_{1:p}D_{1:p}^{-1})-\omega_{\mathrm{qr}}\bigl{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\bigr{\rVert}_{\mathsf{F}},$
$\bigl{\|}\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\bigr{\rVert}_{\mathsf{F}}$ $\leq(1+\omega_{\mathrm{qr}})\bigl{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\bigr{\rVert}_{\mathsf{F}},$
which proves the result under the assumptions (13) and (14).
Then we consider the case when $\bar{V}_{1:p+1}$ is close orthonormal instead of being exactly orthonormal, i.e, $20<\bigl{\|}\bar{V}_{1:p}^{\top}\bar{V}_{1:p}-I\bigr{\rVert}_{\mathsf{F}}\leq\omega_{p}\leq 1/$ 2. In this case, we need to avoid to directly considering $\bar{V}_{1:p+1}$ since $\bar{V}_{1:p+1}$ may be not well-conditioned. Observe that
$\bar{V}_{1:p+1}\widehat{R}_{1:p+1,2:p+1}=\bar{V}_{1:p}\widehat{R}_{1:p,2:p+1}+\bar{V}_{p+1}\widehat{R}_{p+1,2:p+1}$
and further
$\begin{split}\bar{V}_{1:p+1}&\begin{bmatrix}\phi\widehat{R}_{1:p+1,1}&\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\\ &=\bar{V}_{1:p}\begin{bmatrix}\phi\widehat{R}_{1:p,1}&\widehat{R}_{1:p,2:p+1}D_{1:p}^{-1}\end{bmatrix}+\bar{V}_{p+1}\widehat{R}_{p+1,1:p+1}D_{1:p}^{-1}.\end{split}$
From (9), we obtain
$\begin{split}[\phi r,\widehat{W}_{1:p}D_{1:p}^{-1}]+\Delta F=\bar{V}_{1:p}\begin{bmatrix}\phi\widehat{R}_{1:p,1}&\widehat{R}_{1:p,2:p+1}D_{1:p}^{-1}\end{bmatrix}\end{split}$ (29)
with $\Delta F=\Delta E_{1}(\phi)+\Delta E_{1:p+1}(\phi)D_{1:p}^{-1}-\bar{V}_{p+1}\widehat{R}_{p+1,1:p+1}D_{1:p}^{-1}$ satisfying
$\bigl{\|}\Delta F\bigr{\rVert}_{\mathsf{F}}\leq\omega_{\mathrm{qr}}\big{\|}[\phi r,\widehat{W}_{1:p}D_{1:p}^{-1}]\big{\rVert}_{\mathsf{F}}+\omega_{H}\bigl{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\bigr{\rVert}_{\mathsf{F}},$
which also implies, by (13) and Lemma 2,
$\begin{split}&\big{\|}\begin{bmatrix}\phi\widehat{R}_{1:p,1}&\widehat{R}_{1:p,2:p+1}D_{1:p}^{-1}]\end{bmatrix}\big{\rVert}\\ &\qquad\leq\frac{1+\omega_{p}}{1-\omega_{p}}\bigl{(}(1+\omega_{\mathrm{qr}})\big{\|}[\phi r,\widehat{W}_{1:p}D_{1:p}^{-1}]\big{\rVert}_{\mathsf{F}}+\omega_{H}\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}\bigr{)}\\ &\qquad\leq 3\bigl{(}(1+\omega_{\mathrm{qr}})\big{\|}[\phi r,\widehat{W}_{1:p}D_{1:p}^{-1}]\big{\rVert}_{\mathsf{F}}+\omega_{H}\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}\bigr{)}\end{split}$ (30)
and
$\begin{split}\sigma_{\min}\big{(}[\phi\widehat{R}_{1:p,1}&\quad\widehat{R}_{1:p,2:p+1}D_{1:p}^{-1}]\big{)}\\ \leq{}&3\bigl{(}(\omega_{\kappa(r,W)}+\omega_{\mathrm{qr}})\big{\|}[\phi r,\widehat{W}_{1:p}D_{1:p}^{-1}]\big{\rVert}_{\mathsf{F}}+\omega_{H}\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}\bigr{)}.\end{split}$ (31)
Together with
$\big{\|}[\phi\widehat{R}_{1:p+1,1}\quad\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}]\big{\rVert}\leq\big{\|}[\phi\widehat{R}_{1:p,1}\quad\widehat{R}_{1:p,2:p+1}D_{1:p}^{-1}]\big{\rVert}+\bigl{\lvert}\widehat{R}_{p+1,p+1}D_{p,p}^{-1}\bigr{\rvert}$
and
$\begin{split}&\sigma_{\min}\big{(}[\phi\widehat{R}_{1:p+1,1}\quad\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}]\big{)}\\ &\qquad\leq\sigma_{\min}\big{(}[\phi\widehat{R}_{1:p,1}\quad\widehat{R}_{1:p,2:p+1}D_{1:p}^{-1}]\big{)}+\bigl{\lvert}\widehat{R}_{p+1,p+1}D_{p,p}^{-1}\bigr{\rvert},\end{split}$
we prove (18) and (19).
It remains to prove (20) and (21). From Lemma 2 and
$\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}=\widehat{R}_{1:p,2:p+1}D_{1:p}^{-1}+\widehat{R}_{p+1,2:p+1}D_{1:p}^{-1},$
$\widehat{W}_{1:p}D_{1:p}^{-1}+\Delta E_{2:p+1}(\phi)D_{1:p}^{-1}-\bar{V}_{p+1}\widehat{R}_{p+1,2:p+1}D_{1:p}^{-1}=\bar{V}_{1:p}\widehat{R}_{1:p,2:p+1}D_{1:p}^{-1},$
we have
$\begin{split}\sigma_{\min}(\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1})\geq{}&\sigma_{\min}(\widehat{R}_{1:p,2:p+1}D_{1:p}^{-1})-\bigl{\lvert}\widehat{R}_{p+1,p+1}D_{p,p}^{-1}\bigr{\rvert}\\ \geq{}&\frac{1}{1+\omega_{p}}\bigl{(}\sigma_{\min}(\widehat{W}_{1:p}D_{1:p}^{-1})-\bigl{\|}\Delta E_{2:p+1}(\phi)D_{1:p}^{-1}\bigr{\rVert}_{\mathsf{F}}\\ &-\big{\|}\bar{V}_{p+1}\widehat{R}_{p+1,2:p+1}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}\bigr{)}-\bigl{\lvert}\widehat{R}_{p+1,p+1}D_{p,p}^{-1}\bigr{\rvert}\\ \geq{}&\biggl{(}\frac{2}{3}(\omega_{\kappa(W)}-\omega_{\mathrm{qr}}-\omega_{H})-\omega_{H}\biggr{)}\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}},\end{split}$
and similarly
$\big{\|}\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}\leq\biggl{(}\frac{1+\omega_{p}}{1-\omega_{p}}(1+\omega_{\mathrm{qr}}+\omega_{H})+\omega_{H}\biggr{)}\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}$
which proves (20) and (21). ∎
We are now prepared to bound $\bigg{\|}\widehat{R}_{1:p+1}\begin{bmatrix}1\\ -\widehat{y}^{(k)}\end{bmatrix}\bigg{\rVert}$.
Lemma 3.
Assume that $\widehat{y}^{(k)}$ satisfies (15). If (13) and (14) hold with
$\begin{split}\omega_{\kappa(W)}\geq{}&\frac{27\bigl{(}\omega_{\kappa(r,W)}+\omega_{\mathrm{qr}}+\varepsilon_{ls}\bigr{)}}{1-\varepsilon_{r}-6\cdot(\omega_{\kappa(r,W)}+\omega_{\mathrm{qr}})-9\cdot\varepsilon_{ls}}+9\cdot\omega_{\kappa(r,W)}\\ &+10\cdot\omega_{\mathrm{qr}}+9\cdot\varepsilon_{ls}+16\cdot\omega_{H},\end{split}$ (32)
then
$\begin{split}\bigg{\|}\widehat{R}_{1:p+1}\begin{bmatrix}1\\ -\widehat{y}^{(k)}\end{bmatrix}\bigg{\rVert}\leq{}&(9\cdot\omega_{\kappa(r,W)}+9\cdot\omega_{\mathrm{qr}}+12\cdot\varepsilon_{ls})\beta+(9\cdot\omega_{\kappa(r,W)}+9\cdot\omega_{\mathrm{qr}}\\ &+12\cdot\varepsilon_{ls}+12\cdot\omega_{H})\bigl{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\bigr{\rVert}_{\mathsf{F}}\bigl{\|}D_{1:p}\widehat{y}^{(k)}\bigr{\rVert},\end{split}$ (33)
where $\omega_{H}$ is defined in Lemma 1.
Proof.
In [19, Theorem 2.4], we define $\tilde{R}_{1:p+1}=\widehat{R}_{1:p+1}+\Delta R_{1:p+1}$,
$\phi=\bigl{\|}D_{1:p}\widehat{y}^{(i)}\bigr{\rVert}^{-1}>0,\quad\delta(\phi)=\frac{\sigma_{\min}\big{(}\begin{bmatrix}\phi\tilde{R}_{1:p+1,1}&\tilde{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\big{)}}{\sigma_{\min}(\tilde{R}_{1:p+1,2:p+1}D_{1:p}^{-1})}<1.$ (34)
Then
$\begin{split}\bigg{\|}(\widehat{R}_{1:p+1}&+\Delta R_{1:p+1})\begin{bmatrix}1\\ -\widehat{y}^{(k)}\end{bmatrix}\bigg{\rVert}^{2}\\ \leq{}&\min_{y}\bigg{\|}(\widehat{R}_{1:p+1}+\Delta R_{1:p+1})\begin{bmatrix}1&0\\ 0&D_{1:p}^{-1}\end{bmatrix}\begin{bmatrix}1\\ -D_{1:p}y\end{bmatrix}\bigg{\rVert}^{2}\\ \leq{}&\sigma_{\min}^{2}\big{(}\begin{bmatrix}\phi\tilde{R}_{1:p+1,1}&\tilde{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\big{)}\big{\|}D_{1:p}\widehat{y}^{(i)}\big{\rVert}^{2}\biggl{(}1+\frac{1}{1-\delta^{2}(\phi)}\biggr{)}.\end{split}$ (35)
Then our aim is to bound $\sigma_{\min}\big{(}\begin{bmatrix}\phi\tilde{R}_{1:p+1,1}&\tilde{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\big{)}$ and $\delta(\phi)$. First, we will give the bound for $\sigma_{\min}\big{(}\begin{bmatrix}\phi\tilde{R}_{1:p+1,1}&\tilde{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\big{)}$. By (9), (15), (18), and (19),
$\begin{split}&\sigma_{\min}\big{(}[\phi\tilde{R}_{1:p+1,1}\quad\tilde{R}_{1:p+1,2:p+1}D_{1:p}^{-1}]\big{)}\\ \leq{}&\sigma_{\min}\Big{(}\begin{bmatrix}\phi\widehat{R}_{1:p+1,1}&\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\Big{)}+\Big{\|}\begin{bmatrix}\phi\Delta R_{1:p+1,1}&\Delta R_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\Big{\rVert}_{\mathsf{F}}\\ \leq{}&\sigma_{\min}\Big{(}\begin{bmatrix}\phi\widehat{R}_{1:p+1,1}&\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\Big{)}+\varepsilon_{ls}\Big{\|}\begin{bmatrix}\phi\widehat{R}_{1:p+1,1}&\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\Big{\rVert}_{\mathsf{F}}\\ \leq{}&(\alpha_{11}+\varepsilon_{ls}\alpha_{21})\Big{\|}\begin{bmatrix}\phi r,\widehat{W}_{1:p}D_{1:p}^{-1}\end{bmatrix}\Big{\rVert}_{\mathsf{F}}+(\alpha_{12}+\varepsilon_{ls}\alpha_{22})\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}\\ ={}&(\alpha_{11}+\varepsilon_{ls}\alpha_{21})\phi\beta+(\alpha_{11}+\varepsilon_{ls}\alpha_{21}+\alpha_{12}+\varepsilon_{ls}\alpha_{22})\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}},\end{split}$ (36)
which implies that, by noting that $\phi=\bigl{\|}D_{1:p}\widehat{y}^{(i)}\bigr{\rVert}^{-1}$,
$\begin{split}&\sigma_{\min}\big{(}[\phi\tilde{R}_{1:p+1,1}\quad\tilde{R}_{1:p+1,2:p+1}D_{1:p}^{-1}]\big{)}\bigl{\|}D_{1:p}\widehat{y}^{(i)}\bigr{\rVert}\\ &\quad\leq(\alpha_{11}+\varepsilon_{ls}\alpha_{21})\beta+(\alpha_{11}+\varepsilon_{ls}\alpha_{21}+\alpha_{12}+\varepsilon_{ls}\alpha_{22})\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}\bigl{\|}D_{1:p}\widehat{y}^{(i)}\bigr{\rVert}.\end{split}$ (37)
Observing that
$(\widehat{R}_{1:p+1}+\Delta R_{1:p+1})$ $\begin{bmatrix}1\\ -\widehat{y}^{(k)}\end{bmatrix}$
[…]
from (35), (37), Lemma 1, and (15), assuming $\delta(\phi)\leq 1/2$,
$\begin{split}\phi\beta&\leq\frac{1}{1-\varepsilon_{r}}\biggl{(}\phi\bigg{\|}(\widehat{R}_{1:p+1}+\Delta R_{1:p+1})\begin{bmatrix}1\\ -\widehat{y}^{(k)}\end{bmatrix}\bigg{\rVert}+\big{\|}\widehat{R}_{1:p+1,2:p+1}\widehat{y}^{(k)}\big{\rVert}\\ &\quad+\big{\|}\Delta R_{1:p+1,2:p+1}\widehat{y}^{(k)}\big{\rVert}\biggr{)}\\ &\leq\frac{3(\alpha_{11}+\alpha_{12})+3\cdot\varepsilon_{ls}(\alpha_{21}+\alpha_{22})+(1+\varepsilon_{ls})\alpha_{4}}{1-\varepsilon_{r}}\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}\\ &\quad+\frac{3\cdot\alpha_{11}+3\cdot\varepsilon_{ls}\alpha_{21}}{1-\varepsilon_{r}}\phi\beta.\end{split}$
Thus, we bound $\phi\beta$ by $\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}$ as follows:
$\phi\beta\leq\frac{3(\alpha_{11}+\alpha_{12})+3\cdot\varepsilon_{ls}(\alpha_{21}+\alpha_{22})+(1+\varepsilon_{ls})\alpha_{4}}{1-\varepsilon_{r}-3\cdot\alpha_{11}-3\cdot\varepsilon_{ls}\alpha_{21}}\cdot\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}.$
Together with (36) and Lemma 1, we give the bound on $\delta(\phi)$,
$\begin{split}\delta(\phi)\leq{}&\frac{(\alpha_{11}+\varepsilon_{ls}\alpha_{21})\phi\beta+(\alpha_{11}+\varepsilon_{ls}\alpha_{21}+\alpha_{12}+\varepsilon_{ls}\alpha_{22})\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}}{\alpha_{3}\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}}\\ \leq{}&\frac{(\alpha_{11}+\varepsilon_{ls}\alpha_{21})\frac{2\cdot(\alpha_{11}+\alpha_{12})+3\cdot\varepsilon_{ls}(\alpha_{21}+\alpha_{22})+\alpha_{4}}{1-\varepsilon_{r}-2\cdot\alpha_{11}-3\cdot\varepsilon_{ls}\alpha_{21}}+(\alpha_{11}+\varepsilon_{ls}\alpha_{21}+\alpha_{12}+\varepsilon_{ls}\alpha_{22})}{\alpha_{3}}\\ \leq{}&\frac{1}{2},\end{split}$ (38)
which is guaranteed by the assumption (32) with ignoring the quadratic terms. Then combining (35) with (37) and (38), it follows that
$\begin{split}\bigg{\|}\widehat{R}_{1:p+1}\begin{bmatrix}1\\ -\widehat{y}^{(k)}\end{bmatrix}\bigg{\rVert}\leq{}&\bigg{\|}(\widehat{R}_{1:p+1}+\Delta R_{1:p+1})\begin{bmatrix}1\\ -\widehat{y}^{(k)}\end{bmatrix}\bigg{\rVert}+\bigg{\|}\Delta R_{1:p+1}\begin{bmatrix}1\\ -\widehat{y}^{(k)}\end{bmatrix}\bigg{\rVert}\\ \leq{}&(3\cdot\alpha_{11}+4\cdot\varepsilon_{ls}\alpha_{21})\beta+(3\cdot\alpha_{11}+4\cdot\varepsilon_{ls}\alpha_{21}+3\cdot\alpha_{12}\\ &+4\cdot\varepsilon_{ls}\alpha_{22})\big{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\big{\rVert}_{\mathsf{F}}\bigl{\|}D_{1:p}\widehat{y}^{(k)}\bigr{\rVert},\end{split}$
which concludes the proof by substituting $\alpha_{*}$ from Lemma 1 and ignoring the quadratic terms. ∎
The following theorem gives the upper bound on $\bigl{\|}b-A\widehat{x}^{(k)}\bigr{\rVert}$.
Theorem 1.
Assume that $\widehat{W}_{1:is}$, $\widehat{V}_{is+1}$, $\widehat{R}_{is+1,is+1}$, $\widehat{G}_{1:is,1:is}$, $\widehat{T}_{1:is,1:is}$, $\widehat{y}^{(i)}$, and $\widehat{x}^{(i)}$ satisfy (5)–(9), (15), and (17). If there exists an iteration $p:=i^{*}s+j^{*}\leq n$ such that (13) and
$\omega_{\kappa(W)}\frac{\sqrt{p}\,(1+\varepsilon_{AZ}+\delta_{M_{L}^{-1}AZ})\kappa(M_{L})\kappa(A)\kappa(\tilde{Z}_{1:p})}{1-\sqrt{p}\,(\varepsilon_{AZ}+\delta_{M_{L}^{-1}AZ})\kappa(M_{L})\kappa(A)\kappa(\tilde{Z}_{1:p})}\leq 1$ (39)
hold with (32), then
$\begin{split}\bigl{\|}b-A\widehat{x}^{(i^{*})}\bigr{\rVert}&\leq\xi_{b}\kappa(M_{L})\bigl{\|}b\bigr{\rVert}+\xi_{Ax}\frac{\sqrt{p}\,\kappa(\tilde{Z}_{1:p})}{1-\sqrt{p}\,\varepsilon_{mv}\kappa(\tilde{Z}_{1:p})}\kappa(M_{L})\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\bigl{\|}\widehat{x}^{(i^{*})}\bigr{\rVert}\end{split}$ (40)
with
$\xi_{b}={}$ $\sqrt{1+\omega_{p+1}}\,(9\cdot\omega_{\kappa(r,W)}+10\cdot\omega_{\mathrm{qr}}+12\cdot\varepsilon_{ls})(1+2\cdot\delta_{M_{L}^{-1}b}),$
$\xi_{Ax}={}$ $\sqrt{1+\omega_{p+1}}\,(1+\varepsilon_{AZ}+\delta_{M_{L}^{-1}AZ})(9\cdot\omega_{\kappa(r,W)}+10\cdot\omega_{\mathrm{qr}}+12\cdot\varepsilon_{ls}+12\cdot\omega_{H})$
$+\varepsilon_{AZ}+\delta_{M_{L}^{-1}AZ}+\varepsilon_{mv},$
where $\widehat{Z}_{1:p}=\tilde{Z}_{1:p}D_{1:p}$ with a positive definite diagonal matrix $D_{1:p}$, and
$\kappa(\tilde{Z}_{1:p})\leq\frac{(1+\sqrt{p}\,\delta_{M_{R}^{-1}})\kappa(M_{R})\kappa(\tilde{B}_{1:p})}{1-\sqrt{p}\,\delta_{M_{R}^{-1}}\kappa(M_{R})\kappa(\tilde{B}_{1:p})}$
with $\widehat{B}_{1:p}=\tilde{B}_{1:p}D_{1:p}$.
Proof.
We consider the $i^{*}$-th block iteration. From (5), (9), (11), and (12), we have
$\begin{bmatrix}\widehat{r}&M_{L}^{-1}A\widehat{Z}_{1:p}\end{bmatrix}$ $=\begin{bmatrix}\widehat{r}&\widehat{W}_{1:p}\end{bmatrix}-\begin{bmatrix}0&\Delta W_{1:p}\end{bmatrix}$
$=\bar{V}_{1:p+1}\widehat{R}_{1:p+1}(1)-\underbrace{\bigl{(}\Delta E_{1:p+1}(1)+\begin{bmatrix}0&\Delta W_{1:p}\end{bmatrix}\bigr{)}}_{=:\Delta F_{1:p+1}}$
with
$\begin{split}\bigg{\|}\Delta F_{1:p+1}\begin{bmatrix}1\\ -\widehat{y}^{(i^{*})}\end{bmatrix}\bigg{\rVert}\leq{}&\bigg{\|}\Delta E_{1:p+1}(1)\begin{bmatrix}1\\ -\widehat{y}^{(i^{*})}\end{bmatrix}\bigg{\rVert}+\bigg{\|}\begin{bmatrix}0&\Delta W_{1:p}\end{bmatrix}\begin{bmatrix}1\\ -\widehat{y}^{(i^{*})}\end{bmatrix}\bigg{\rVert}\\ \leq{}&\omega_{\mathrm{qr}}\beta+\bigl{(}\omega_{\mathrm{qr}}\bigl{(}1+\varepsilon_{AZ}+\delta_{M_{L}^{-1}AZ}\bigr{)}+\varepsilon_{AZ}+\delta_{M_{L}^{-1}AZ}\bigr{)}\\ &\cdot\bigl{\|}M_{L}^{-1}\bigr{\rVert}\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\bigl{\|}\tilde{Z}_{1:p}\bigr{\rVert}_{\mathsf{F}}\bigl{\|}D_{1:p}\widehat{y}^{(i^{*})}\bigr{\rVert}.\end{split}$ (41)
Together with (5), (17), and Lemma 3, we have
$\begin{split}&\bigl{\|}M_{L}^{-1}(b-A\widehat{x}^{(i^{*})})\bigr{\rVert}\\ &\quad\leq\bigl{\|}M_{L}^{-1}b-M_{L}^{-1}A\widehat{Z}_{1:p}\widehat{y}^{(i^{*})}\bigr{\rVert}+\bigl{\|}M_{L}^{-1}A\Delta x^{(i^{*})}\bigr{\rVert}\\ &\quad\leq\bigg{\|}\begin{bmatrix}\widehat{r}&M_{L}^{-1}A\widehat{Z}_{1:p}\end{bmatrix}\begin{bmatrix}1\\ -\widehat{y}^{(i^{*})}\end{bmatrix}\bigg{\rVert}+\bigl{\|}\Delta r\bigr{\rVert}+\bigl{\|}M_{L}^{-1}\bigr{\rVert}\bigl{\|}A\bigr{\rVert}\bigl{\|}\Delta x^{(i^{*})}\bigr{\rVert}\\ &\quad\leq\bigg{\|}\bar{V}_{1:p+1}\widehat{R}_{1:p+1}(1)\begin{bmatrix}1\\ -\widehat{y}^{(i^{*})}\end{bmatrix}\bigg{\rVert}+\bigg{\|}\Delta F_{1:p+1}\begin{bmatrix}1\\ -\widehat{y}^{(i^{*})}\end{bmatrix}\bigg{\rVert}+\bigl{\|}\Delta r\bigr{\rVert}\\ &\qquad+\bigl{\|}M_{L}^{-1}\bigr{\rVert}\bigl{\|}A\bigr{\rVert}\bigl{\|}\Delta x^{(i^{*})}\bigr{\rVert}\\ &\quad\leq\sqrt{1+\omega_{p+1}}\,\bigg{\|}\widehat{R}_{1:p+1}(1)\begin{bmatrix}1\\ -\widehat{y}^{(i^{*})}\end{bmatrix}\bigg{\rVert}+\bigg{\|}\Delta F_{1:p+1}\begin{bmatrix}1\\ -\widehat{y}^{(i^{*})}\end{bmatrix}\bigg{\rVert}\\ &\qquad+\delta_{M_{L}^{-1}b}\bigl{\|}M_{L}^{-1}\bigr{\rVert}\bigl{\|}b\bigr{\rVert}+\varepsilon_{mv}\bigl{\|}M_{L}^{-1}\bigr{\rVert}\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\bigl{\|}\tilde{Z}_{1:p}\bigr{\rVert}_{\mathsf{F}}\bigl{\|}D_{1:p}\widehat{y}^{(i^{*})}\bigr{\rVert}\\ &\quad\leq\sqrt{1+\omega_{p+1}}\,(9\cdot\omega_{\kappa(r,W)}+10\cdot\omega_{\mathrm{qr}}+12\cdot\varepsilon_{ls})(1+2\cdot\delta_{M_{L}^{-1}b})\bigl{\|}M_{L}^{-1}\bigr{\rVert}\bigl{\|}b\bigr{\rVert}\\ &\qquad+\bigl{(}\sqrt{1+\omega_{p+1}}\,(1+\varepsilon_{AZ}+\delta_{M_{L}^{-1}AZ})\cdot(9\cdot\omega_{\kappa(r,W)}+10\cdot\omega_{\mathrm{qr}}+12\cdot\varepsilon_{ls}\\ &\qquad+12\cdot\omega_{H})+\varepsilon_{AZ}+\delta_{M_{L}^{-1}AZ}+\varepsilon_{mv}\bigr{)}\bigl{\|}M_{L}^{-1}\bigr{\rVert}\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\bigl{\|}\tilde{Z}_{1:p}\bigr{\rVert}_{\mathsf{F}}\bigl{\|}D_{1:p}\widehat{y}^{(i^{*})}\bigr{\rVert}.\end{split}$ (42)
Note that using Lemma 3 requires (14) which is guaranteed by the assumption (39).
Then it remains to bound $\bigl{\|}D_{1:p}\widehat{y}^{(i^{*})}\bigr{\rVert}$. By
$\begin{split}\sigma_{\min}(\tilde{Z}_{1:p})\bigl{\|}D_{1:p}\widehat{y}^{(i^{*})}\bigr{\rVert}&\leq\bigl{\|}\tilde{Z}_{1:p}(D_{1:p}\widehat{y}^{(i^{*})})\bigr{\rVert}\\ &\leq\bigl{\|}x^{(i^{*})}\bigr{\rVert}+\bigl{\|}\Delta x^{(i^{*})}\bigr{\rVert}\\ &\leq\bigl{\|}x^{(i^{*})}\bigr{\rVert}+\varepsilon_{mv}\bigl{\|}\tilde{Z}_{1:p}\bigr{\rVert}_{\mathsf{F}}\bigl{\|}D_{1:p}\widehat{y}^{(i^{*})})\bigr{\rVert},\end{split}$
we then derive
$\bigl{\|}D_{1:p}\widehat{y}^{(i^{*})}\bigr{\rVert}\leq\frac{\bigl{\|}\widehat{x}^{(i^{*})}\bigr{\rVert}}{\sigma_{\min}(\tilde{Z}_{1:p})-\varepsilon_{mv}\bigl{\|}\tilde{Z}_{1:p}\bigr{\rVert}_{\mathsf{F}}}=\frac{\bigl{\|}\widehat{x}^{(i^{*})}\bigr{\rVert}}{\sigma_{\min}(\tilde{Z}_{1:p})\bigl{(}1-\sqrt{p}\,\varepsilon_{mv}\kappa(\tilde{Z}_{1:p})\bigr{)}},$
which proves the conclusion combined with (42) and $\bigl{\|}b-A\widehat{x}^{(i^{*})}\bigr{\rVert}\leq\bigl{\|}M_{L}\bigr{\rVert}\bigl{\|}M_{L}^{-1}(b-A\widehat{x}^{(i^{*})})\bigr{\rVert}$. ∎
For simplicity, Theorem 1 does not consider the restarted $s$-step GMRES algorithm. Analogous to [5, Theorem 4.1], it is easy to generalize Theorem 1 to the restarted $s$-step GMRES algorithm, which can be regarded as an iterative refinement to remove $\kappa(M_{L})$, $\kappa(M_{R})$, and $\kappa(\tilde{B}_{1:p})$ from the bound of the backward error in Theorem 1 under certain conditions.
Remark 2.
According to Remark 1, $\varepsilon_{r}$, $\varepsilon_{AZ}$, $\varepsilon_{ls}$, $\varepsilon_{mv}=O(\bm{u})$, and $\delta_{*}$ are highly preconditioner dependent. It remains to evaluate $1+\omega_{p+1}$, $\omega_{\mathrm{qr}}$, $\omega_{\kappa(r,W)}$, and $\omega_{H}$, which are all related only to the block orthogonalization method. Among these terms, it is usually easy to check that $1+\omega_{p+1}\leq O(\sqrt{n}\,)$ and $\omega_{\mathrm{qr}}=O(\bm{u})$ for different block orthogonalization methods.
Then from Theorem 1 the difficulty of analyzing the backward error of the $s$-step GMRES algorithm is to prove that (13) is satisfied when $\widehat{V}_{1:p+1}$ is not well-conditioned, namely, to determine $\omega_{\kappa(r,W)}$. In the case that $\bar{V}_{*}$ is not exactly orthonormal, we need to additionally estimate $\omega_{H}$.
We now turn to the analysis of the backward stability of $s$-step GMRES with specific block orthogonalization schemes. As we will see for all orthogonalization methods in Section 4, there is a dependence of the backward error on $\kappa(\tilde{B}_{i:p})$. It is not clear how to bound this quantity a priori. We will give an example in Section 6 that shows that $\kappa(\tilde{B}_{i:p})$ can be large even when the condition numbers of its blocks are small, and then we will present a modified approach to eliminate this problem.
4 Backward stability of $s$-step GMRES with different orthogonalization methods
Based on the framework introduced in Section 3, we now analyze the $s$-step GMRES algorithm combined with specific block orthogonalization methods. Note that for simplicity, we assume $M_{L}=I$ and $M_{R}=I$ in this section, because the constants related to preconditioners vary significantly depending on the particular preconditioners chosen. Consequently, $\kappa(M_{L})=\kappa(M_{R})=1$, $\kappa(\tilde{Z})=\kappa(\tilde{B})$, and $\delta_{M_{L}^{-1}b}=\delta_{M_{L}^{-1}AZ}=0$. For readers who wish to analyze the backward error when using specific preconditioners, only $\kappa(M_{L})$, $\kappa(\tilde{Z})$, $\delta_{M_{L}^{-1}b}$, and $\delta_{M_{L}^{-1}AZ}$ in Theorem 1 must be determined.
The bounds for the relative backward error derived in this section will all ultimately depend on the condition number of the computed basis for the Krylov subspace. This confirms what has been widely observed experimentally, i.e., that the conditioning of the Krylov basis is crucial in determining the resulting numerical behavior of $s$-step GMRES. It is, unfortunately, difficult to bound this quantity a priori. In Section 6, we comment on attempting to control the condition number and present a modified approach for ensuring that this quantity remains small.
4.1 Backward error of $s$-step GMRES with block Householder QR and block modified Gram–Schmidt (BMGS)
In [21], Walker discussed the GMRES algorithm with Householder QR orthogonalization, which can be directly extended to $s$-step GMRES with block Householder QR. Note that using Householder QR in GMRES is more computationally intensive than alternative orthogonalization schemes. As a result, there is limited research on employing block Householder QR in the $s$-step GMRES algorithm. However, for purposes of illustration, we also analyze the backward stability of this variant.
From [13, Theorem 19.4] and [23], it is easy to see that
$\begin{bmatrix}\phi\widehat{r},\widehat{W}_{1:n-1}\end{bmatrix}+\Delta E_{1:n}(\phi)=\tilde{V}_{1:n}\widehat{R}_{1:n}(\phi),\quad\bigl{\|}\Delta E_{j}(\phi)\bigr{\rVert}\leq O(\bm{u})\Big{\|}\begin{bmatrix}\widehat{r},\widehat{W}_{1:n-1}\end{bmatrix}_{j}\Big{\rVert},$ (43)
for any $j\in\{1,\dotsc,n\}$, is satisfied with an exact orthogonal matrix $\tilde{V}_{1:n}$ for the block Householder QR algorithm. This means that for block Householder QR orthogonalization, (13) is satisfied with $\omega_{\kappa(r,W)}=O(\bm{u})$ when $p=n$, since the block Householder QR algorithm generates a well-conditioned set of vectors until $p=n$.
Then we consider the modified Gram–Schmidt (MGS) algorithm. [13, Theorem 19.13] and [4] showed (43) can be satisfied by establishing the equivalence between the MGS algorithm and the Householder QR algorithm. Similarly, it has been proved in [2] that applying the BMGS algorithm with Householder QR as the intra-block orthogonalization routine to $X$ is equivalent to applying the block Householder QR algorithm to $\begin{bmatrix}0_{n}\\ X\end{bmatrix}$ both mathematically and numerically. Then we obtain that (43) holds also for the BMGS algorithm.
Unlike the block Householder QR algorithm, we cannot ensure the orthogonality of $\widehat{V}_{1:n}$. Therefore, it is necessary to demonstrate that $\bigl{\|}\widehat{V}_{1:j}^{\top}\widehat{V}_{1:j}-I\bigr{\rVert}_{\mathsf{F}}\leq 1/2$ if
$\sigma_{\min}\Big{(}\begin{bmatrix}\phi\widehat{r},\widehat{W}_{1:j}D_{1:j}^{-1}\end{bmatrix}\Big{)}/2\geq O(\bm{u})\Big{\|}\begin{bmatrix}\phi\widehat{r},\widehat{W}_{1:j}D_{1:j}^{-1}\end{bmatrix}\Big{\rVert}_{\mathsf{F}}$
holds, as established in [16, Theorem 4.1]. Moreover, the contrapositive indicates that (13) is satisfied with $\omega_{\kappa(r,W)}=O(\bm{u})$ when there exists $p\leq n$ such that $\bigl{\|}\widehat{V}_{1:p+1}^{\top}\widehat{V}_{1:p+1}-I\bigr{\rVert}_{\mathsf{F}}>1/2$ but $\bigl{\|}\widehat{V}_{1:p}^{\top}\widehat{V}_{1:p}-I\bigr{\rVert}_{\mathsf{F}}\leq 1/2$.
Therefore, combined with $\varepsilon_{r}$, $\omega_{\mathrm{qr}}$, $\varepsilon_{AZ}$, $\varepsilon_{ls}$, $\varepsilon_{mv}=O(\bm{u})$, and $\omega_{p}=0$ from Remark 1, Theorem 1 implies the following lemma.
Lemma 4.
Assume that $\widehat{W}_{1:is}$, $\widehat{V}_{is+1}$, $\widehat{R}_{is+1,is+1}$, $\widehat{G}_{1:is,1:is}$, $\widehat{T}_{1:is,1:is}$, $\widehat{y}^{(i)}$, and $\widehat{x}^{(i)}$ satisfying (5)–(9), (15), and (17), are computed by the $s$-step GMRES algorithm with block Householder QR or BMGS orthogonalization. There exists $p=i^{*}s+j^{*}$ such that (13) holds. If it also holds that
$\frac{O(\bm{u})\kappa(A)\kappa(\tilde{B}_{1:p})}{1-O(\bm{u})\kappa(A)\kappa(\tilde{B}_{1:p})}\leq 1,$
then $\kappa(\widehat{V}_{1:p})\leq 3$ and
$\begin{split}\frac{\bigl{\|}b-A\widehat{x}^{(i^{*})}\bigr{\rVert}}{\bigl{\|}b\bigr{\rVert}+\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\bigl{\|}\widehat{x}^{(i^{*})}\bigr{\rVert}}\leq{}&\frac{O(\bm{u})\kappa(\tilde{B}_{1:p})}{1-O(\bm{u})\kappa(\tilde{B}_{1:p})},\end{split}$
where $\widehat{B}_{1:p}=\tilde{B}_{1:p}D_{1:p}$ with a positive definite diagonal matrix $D_{1:p}$.
Furthermore, if $s=1$, as long as $\frac{O(\bm{u})\kappa(A)}{1-O(\bm{u})\kappa(A)}\leq 1,$ then there exists $p=i^{*}s+j^{*}$ such that $\kappa(\widehat{B}_{1:p})=\kappa(\widehat{V}_{1:p})\leq 3$ and
$\begin{split}\frac{\bigl{\|}b-A\widehat{x}^{(i^{*})}\bigr{\rVert}}{\bigl{\|}b\bigr{\rVert}+\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\bigl{\|}\widehat{x}^{(i^{*})}\bigr{\rVert}}\leq{}&\frac{O(\bm{u})}{1-O(\bm{u})}.\end{split}$
For $s=1$, i.e., the standard GMRES algorithm with Householder QR and MGS orthogonalization, we recover the same backward stability result of [12] and [19], respectively.
4.2 Backward error of $s$-step GMRES with reorthogonalized block classical Gram–Schmidt (BCGSI+)
In this subsection, we consider a different and more complicated situation, for example BCGSI+, which means that we cannot obtain (43) with an exactly orthonormal matrix $\tilde{V}$. According to Lemma 3, we need to prove both (13) and (26). We provide the properties of BCGSI+ in Appendix A.
By Lemma 7, we obtain that $\omega_{\mathrm{qr}}=O(\bm{u})$ and (13) holds with $\omega_{\kappa(r,W)}=O(\bm{u})$. From Lemma 8 and the fact that $\widehat{T}$ is upper triangular, (26) holds with $\omega_{H}=O(\bm{u})$. Together with $\varepsilon_{r}$, $\varepsilon_{AZ}$, $\varepsilon_{ls}$, $\varepsilon_{mv}=O(\bm{u})$ as discussed in Remark 1, we derive the following lemma to show the backward stability of the $s$-step GMRES with the BCGSI+ algorithm using Theorem 1.
Lemma 5.
Assume that $\widehat{W}_{1:is}$, $\widehat{V}_{is+1}$, $\widehat{R}_{is+1,is+1}$, $\widehat{G}_{1:is,1:is}$, $\widehat{T}_{1:is,1:is}$, $\widehat{y}^{(i)}$, and $\widehat{x}^{(i)}$ satisfying (5)–(9), (15), and (17), are computed by the $s$-step GMRES algorithm with BCGSI+ orthogonalization. There exists $p=i^{*}s+j^{*}$ such that (13) holds. If it also holds that
$\frac{O(\bm{u})\kappa(A)\kappa(\tilde{B}_{1:p})}{1-O(\bm{u})\kappa(A)\kappa(\tilde{B}_{1:p})}\leq 1,$
then $\kappa(\widehat{V}_{1:p})\leq\frac{1+O(\bm{u})}{1-O(\bm{u})}$ and
$\begin{split}\frac{\bigl{\|}b-A\widehat{x}^{(i^{*})}\bigr{\rVert}}{\bigl{\|}b\bigr{\rVert}+\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\bigl{\|}\widehat{x}^{(i^{*})}\bigr{\rVert}}\leq{}&\frac{O(\bm{u})\kappa(\tilde{B}_{1:p})}{1-O(\bm{u})\kappa(\tilde{B}_{1:p})},\end{split}$
where $\widehat{B}_{1:p}=\tilde{B}_{1:p}D_{1:p}$ with a positive definite diagonal matrix $D_{1:p}$.
Furthermore, if $s=1$, as long as $\frac{O(\bm{u})\kappa(A)}{1-O(\bm{u})\kappa(A)}\leq 1,$ then there exists $p=i^{*}s+j^{*}$ such that $\kappa(\widehat{B}_{1:p})=\kappa(\widehat{V}_{1:p})\leq\frac{1+O(\bm{u})}{1-O(\bm{u})}$ and
$\begin{split}\frac{\bigl{\|}b-A\widehat{x}^{(i^{*})}\bigr{\rVert}}{\bigl{\|}b\bigr{\rVert}+\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\bigl{\|}\widehat{x}^{(i^{*})}\bigr{\rVert}}\leq{}&\frac{O(\bm{u})}{1-O(\bm{u})}.\end{split}$
For $s=1$, i.e., the standard GMRES algorithm with CGSI+ orthogonalization, we recover the same backward stability result of [12].
5 Discussion of theoretical results
In this part, we discuss the stopping criteria and the requirements of the orthogonalization method indicated by the above theoretical results.
5.1 Stopping criteria
The commonly used stopping criteria for the GMRES algorithm are
$\widehat{\beta}\widehat{G}_{1,is+1}\leq\texttt{tolLS}\cdot\bigl{\|}r\bigr{\rVert}$ (44)
and
$\bigl{\|}b-A\widehat{x}^{(i)}\bigr{\rVert}\leq\texttt{tol}\cdot\bigl{(}\bigl{\|}b\bigr{\rVert}+\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\bigl{\|}\widehat{x}^{(i)}\bigr{\rVert}\bigr{)},$ (45)
where tolLS and tol are user-specified thresholds. The first criterion (44) is straightforward and economical to check practically, yet it might fail to recognize timely convergence. This is because $\widehat{\beta}\widehat{G}_{1,is+1}/\bigl{\|}r\bigr{\rVert}$ might not be sufficiently small, even when $\bigl{\|}b-A\widehat{x}^{(i)}\bigr{\rVert}/\bigl{(}\bigl{\|}b\bigr{\rVert}+\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\bigl{\|}\widehat{x}^{(i)}\bigr{\rVert}\bigr{)}$ is sufficiently small. The disadvantage of (45) is that it requires more computational effort, as $\widehat{x}^{(i)}$ and $A\widehat{x}^{(i)}$ must be determined. Consequently, in practice, both criteria are generally employed. This implies that  (44) is checked in every iteration, whereas  (45) is evaluated in select iterations when implementing the GMRES algorithm.
For the $s$-step GMRES with $s=1$, our theoretical results, Lemmas 4 and 5, show that the backward error is $O(\bm{u})$, provided that $A$ is numerically nonsingular. This aligns with the existing results for standard GMRES. However, for $s>1$, the situation differs. From Theorem 1, as well as Lemmas 4 and 5, we find that there is no guarantee for the backward error of the $s$-step GMRES algorithm with $s>1$ due to the existence of $\kappa(\tilde{B}_{1:p})$. This implies that for the “key dimension” $p$, the $s$-step GMRES algorithm nearly reaches optimal accuracy when (13) is satisfied. Meanwhile, if (14) also holds, the backward error can be bounded by $O(\bm{u})\kappa(\tilde{B}_{1:p})$. Otherwise, there is no theoretical guarantee for this “optimal” accuracy. This means that it is not possible to find this “optimal” accuracy only through detecting (44) and (45).
Based on our analysis, the time when $\widehat{V}_{1:p+1}$ loses orthogonality also indicates that
$\frac{\max\bigl{(}\bigl{\lvert}\widehat{R}_{p+1,p+1}D_{p,p}^{-1}\bigr{\rvert},\bigl{\|}\bar{V}_{p+1}\widehat{R}_{p+1,1:p+1}D_{1:p}^{-1}\bigr{\rVert}_{\mathsf{F}}\bigr{)}}{\bigl{\|}\widehat{W}_{1:p}D_{1:p}^{-1}\bigr{\rVert}_{\mathsf{F}}}$
is sufficiently small, as described in (26). Therefore, we introduce an additional criterion to identify the “key dimension” by
$\bigl{\lvert}\widehat{R}_{p+1,p+1}\bigr{\rvert}\leq\texttt{tolH}\cdot\bigl{\|}\widehat{W}_{1:p}\bigr{\rVert}_{\mathsf{F}},$ (46)
for a user-specified threshold tolH, which can terminate the algorithm when it achieves approximately “optimal” accuracy. In standard GMRES, (46) is rarely used, since (45) or (44) is satisfied but (46) is not for many cases. The reason for this is that (46) indicates the quality of the approximation of $A^{-1}$, which is a challenge to estimate.
5.2 Requirement of the orthogonalization method
In Theorem 1 and Remark 2, we show that $\omega_{\kappa(r,W)}$ is determined by the orthogonalization method and has a direct effect on the backward error (40). Revisiting the above proof, we bound $\omega_{\kappa(r,W)}$ by analyzing the loss of orthogonality of the orthogonalization method used for the QR factorization of $X$, specifically $X=QR$. If the orthogonalization method satisfies
$\bigl{\|}\widehat{Q}^{\top}\widehat{Q}-I\bigr{\rVert}_{\mathsf{F}}\leq\frac{1}{2},$ (47)
provided $O(\bm{u})\kappa^{\alpha}(X)\leq 1$, then the contrapositive indicates that $O(\bm{u})\kappa^{\alpha}(X)>1$ when $\widehat{Q}$ is not well-conditioned. Note that when $\widehat{Q}$ is not well-conditioned, it implies that GMRES reaches the key dimension $p$. As mentioned in Section 3, $\omega_{\kappa(r,W)}$, as defined in (13), can be deduced from $O(\bm{u})\kappa^{\alpha}(X)>1$, which amounts to $\sigma_{\min}(X)<O(\bm{u}^{1/\alpha})\bigl{\|}X\bigr{\rVert}_{\mathsf{F}}$. By substituting $\begin{bmatrix}\widehat{r}&\widehat{W}_{1:p}\end{bmatrix}$ for $X$, it follows that $\omega_{\kappa(r,W)}=O(\bm{u}^{1/\alpha})$ in (13).
Furthermore, even for standard GMRES, i.e., $s=1$, employing an orthogonalization method with $\alpha=2$, such as the reorthogonalized Pythagorean variants of BCGS introduced by [8, 10], implies that $\omega_{\kappa(r,W)}=O(\sqrt{\bm{u}}\,)$. Note that other terms, $\omega_{*}$ and $\varepsilon_{*}$, are usually $O(\bm{u})$, except for $\omega_{H}$ which also depends on the orthogonalization method. As a result, ignoring the effect of preconditioning, $\xi_{b}$ and $\xi_{Ax}$ from (40) are dominated by $\omega_{\kappa(r,W)}$, meaning that $\xi_{b}$ and $\xi_{Ax}$ are at least $O(\sqrt{\bm{u}}\,)$. Thus, plugging them into (40), we can only expect the backward error to be bounded by
$\begin{split}\frac{\bigl{\|}b-A\widehat{x}^{(i^{*})}\bigr{\rVert}}{\bigl{\|}b\bigr{\rVert}+\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\bigl{\|}\widehat{x}^{(i^{*})}\bigr{\rVert}}\leq{}&O(\sqrt{\bm{u}}\,),\end{split}$
which is illustrated by an example in Appendix B.
This suggests that we cannot prove a $O(\bm{u})$-level backward error for GMRES or $s$-step GMRES with orthogonalization methods that require $O(\bm{u})\kappa^{\alpha}(X)$ with $\alpha>1$.
6 A modified Arnoldi process for improving the stability of $s$-step GMRES
As indicated in Section 3, the backward stability of $s$-step GMRES algorithm is directly influenced by the condition number of
$\begin{split}\tilde{B}_{1:p}&=\hat{B}_{1:p}D^{-1}_{1:p}\\ &=\begin{bmatrix}\hat{B}_{1:s}D^{-1}_{1:s}&\dotsb&\hat{B}_{(i^{*}-1)s+1:i^{*}s}D^{-1}_{(i^{*}-1)s+1:i^{*}s}&\hat{B}_{i^{*}s+1:p}D^{-1}_{i^{*}s+1:p}\end{bmatrix}\\ &=\begin{bmatrix}\tilde{B}_{1:s}&\dotsb&\tilde{B}_{(i^{*}-1)s+1:i^{*}s}&\tilde{B}_{i^{*}s+1:p}\end{bmatrix}.\end{split}$
In the classical $s$-step GMRES algorithm, the Krylov submatrices $K_{(i-1)s+1:is}$ are utilized as $B_{(i-1)s+1:is}$, with $K_{(i-1)s+1:is}$ being formed as
$K_{(i-1)s+1:is}=\begin{bmatrix}p_{0}(A)V_{(i-1)s+1}&p_{1}(A)V_{(i-1)s+1}&\dotsi&p_{s-1}(A)V_{(i-1)s+1}\end{bmatrix},$ (48)
[…]
One method to manage $\kappa(\tilde{B}_{1:p})$ is to regulate each sub-block of $\tilde{B}_{1:p}$ by adaptively selecting different $s$ in the algorithm, as advocated in [15]. This approach is effective in many scenarios. However, in certain specific cases, $\kappa(\tilde{B}_{1:p})$ can be very large even if the condition number of each sub-block of $\tilde{B}_{1:p}$ is small. This implies that merely constraining the condition number of each sub-block is insufficient to bound $\kappa(\tilde{B}_{1:p})$. We demonstrate this through an example below.
Example 1.
We construct the linear system $Ax=b$, where $A$ is a $20$-by-$20$ random matrix with $\kappa(A)=10^{5}$ generated using the MATLAB commands rng(1) and gallery(’randsvd’, [20, 20], 1e5, 1). The vector $b$ is selected as the right singular vector corresponding to the fourth largest singular value, and the initial guess $x_{0}$ is the zero vector.
For this specific linear system, we use the restarted $s$-step GMRES with monomial/Newton/Chebyshev basis and BCGSI+. For simplicity, we do not consider the preconditioned version. We consider an extreme scenario where $s$-step GMRES undergoes a restart every $20$ iterations, which corresponds to the dimension of $A$. Then it can be ensured that the condition number of each sub-block $\tilde{B}_{(i-1)s:is}$ is less than $1.0\cdot 10^{5}$ by setting $s=3$. However, $\kappa(\tilde{B}_{1:p})>10^{9}$, and the relative backward error $\frac{\bigl{\|}Ax-b\bigr{\rVert}}{\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\bigl{\|}x\bigr{\rVert}+\bigl{\|}b\bigr{\rVert}}$ of the solution computed by the $s$-step GMRES with monomial/Newton/Chebyshev basis is at best around $10^{-8}$, even when $p=20$, as shown in Figure 1. Note that for $s=3$, the backward error can be refined by the restart process.
Unfortunately, the restart process does not always work, for example, $s=4$ as shown in Figure 1. For this case, the condition number of each sub-block $\tilde{B}_{(i-1)s:is}$ is less than $1.0\cdot 10^{10}$, but the relative backward error is at best around $10^{-5}$ even using restart process. This means that we cannot control the condition number of the entire basis through controlling the condition number of each sub-block for the basis, and cannot use a restart process to recover the backward stability.
Figure 1: The plot for Example 1: From left to right, the plots are of relative backward error, the condition number of the basis $\tilde{B}$, and the condition number of the sub-block for the basis of each iteration, where we normalize each column of $\hat{B}$ as $\tilde{B}$. Each line with different color in the plots denotes $s$-step GMRES using different polynomials, including monomial, Newton, and Chebyshev polynomials, to generate the basis by (48).
Observe that $\hat{B}_{1:p}=\hat{V}_{1:p}$ is evidently well-conditioned when $s=1$, and the issue described above arises only when $s>1$. Therefore, to overcome this problem, we aim to ensure that $B_{1:p}$ is near orthonormal, making $B_{1:p}$ well-conditioned. However, directly computing a QR factorization of $K_{1:p}$ is not feasible because $K_{1:p}$ becomes very ill-conditioned as $s$ increases. Note that $\operatorname{span}(V_{1:j})=\operatorname{span}(B_{1:j})$ as established in Theorem 2, and $V_{1:j}$ remains well-conditioned until convergence. Thus instead, we compute a QR factorization of the matrix $\begin{bmatrix}V_{1:(i-2)s+1:(i-1)s}&K_{(i-1)s+1:is}\end{bmatrix}$ during the $i$-th iteration, based on the orthonormal matrix $V_{1:(i-2)s+1:(i-1)s}$. This implies that $B_{(i-1)s+1:is}$ is selected to be the $Q$-factor from the QR factorization of $(I-V_{1:(i-1)s}V_{1:(i-1)s}^{\top})K_{(i-1)s+1:is}$.
The resulting modified $s$-step Arnoldi process is outlined in Algorithm 3. The distinction between the classical and modified $s$-step Arnoldi algorithms lies in the fact that the modified version employs an extra QR factorization, which can be computed by low-synchronization QR algorithms, to obtain $B_{(i-1)s+1:is}$ in Line 2 of Algorithm 3, rather than directly using $K_{(i-1)s+1:is}$ as $B_{(i-1)s+1:is}$ in Line 1 of Algorithm 1. This modified approach allows for the utilization of a significantly larger $s$, with the trade-off of increasing the computation and communication cost of the QR factorization almost twofold; note that for sufficiently large $s$, this still provides an asymptotic communication savings versus standard GMRES.
0:  A matrix $A\in\mathbb{R}^{n\times n}$, a vector $r$, a block size $s$, a left-preconditioner $M_{L}\in\mathbb{R}^{n\times n}$, a right-preconditioner $M_{R}\in\mathbb{R}^{n\times n}$, the basis $B_{1:(i-1)s}$ and the preconditioned basis $Z_{1:(i-1)s}$ generated by the first $i-1$ classical Arnoldi steps, the matrix $W_{1:(i-1)s}$, the orthonormal matrix $V_{1:(i-1)s+1}$, and the upper triangular matrix $R_{1:(i-1)s+1}$ satisfying $\begin{bmatrix}r&W_{1:(i-1)s}\end{bmatrix}=V_{1:(i-1)s+1}R_{1:(i-1)s+1}$.
0:  The basis $B_{1:is}$, the preconditioned basis $Z_{1:is}$, the matrices $W_{1:is}$, $V_{1:is+1}$, and $R_{1:is+1}$ satisfying $\begin{bmatrix}r&W_{1:is}\end{bmatrix}=V_{1:is+1}R_{1:is+1}$.
1:  $K_{(i-1)s+1:is}\leftarrow\begin{bmatrix}p_{0}(A)V_{(i-1)s+1}&p_{1}(A)V_{(i-1)s+1}&\dotsi&p_{s-1}(A)V_{(i-1)s+1}\end{bmatrix}$.
2:  $B_{(i-1)s+1:is}$ is the $Q$-factor of $(I-V_{1:(i-1)s}V_{1:(i-1)s}^{\top})^{2}K_{(i-1)s+1:is}$ satisfying $(I-V_{1:(i-1)s}V_{1:(i-1)s}^{\top})^{2}K_{(i-1)s+1:is}=B_{(i-1)s+1:is}S_{(i-1)s+1:is}$.
3:  $Z_{(i-1)s+1:is}\leftarrow M_{R}^{-1}B_{(i-1)s+1:is}$.
4:  $W_{(i-1)s+1:is}\leftarrow M_{L}^{-1}AZ_{(i-1)s+1:is}$.
5:  Compute the QR factorization of $\begin{bmatrix}r&W_{1:is}\end{bmatrix}=V_{1:is+1}R_{1:is+1}$ based on $\begin{bmatrix}r&W_{1:(i-1)s}\end{bmatrix}=V_{1:(i-1)s+1}R_{1:(i-1)s+1}$.
Algorithm 3 The $i$-th step of the modified $s$-step Arnoldi process
As the above discussion, we first prove that the space spanned by $B_{1:is}$ is the same as the space spanned by $K_{1:is}$, the Krylov basis, in exact arithmetic.
Theorem 2.
Assume that $K_{1:is}$ is defined by (48), and $B_{1:is}$, $V_{1:is+1}$ are obtained via Algorithm 2 with Algorithm 3. Then
$\operatorname{span}\{B_{1:is}\}=\operatorname{span}\{K_{1:is}\}=\operatorname{span}\{V_{1:is}\}=\operatorname{span}\bigl{\{}\begin{bmatrix}r&AB_{1:is-1}\end{bmatrix}\bigr{\}}.$ (49)
Proof.
We prove this theorem by induction. For the base case, $B_{1:s}$ is the $Q$-factor of $K_{1:s}$, i.e., $K_{1:s}=B_{1:s}S_{1:s}$ and $K_{1:s-1}=B_{1:s-1}S_{1:s-1}$, which amounts to
$\operatorname{span}\{B_{1:s-1}\}=\operatorname{span}\{K_{1:s-1}\},\qquad\operatorname{span}\{B_{1:s}\}=\operatorname{span}\{K_{1:s}\}.$
Together with the definition (48) of $K_{1:s}$, we derive
$\begin{split}\operatorname{span}\{V_{1:s}\}={}&\operatorname{span}\bigl{\{}\begin{bmatrix}r&AB_{1:s-1}\end{bmatrix}\bigr{\}}\\ ={}&\operatorname{span}\bigl{\{}\begin{bmatrix}r&AK_{1:s-1}\end{bmatrix}\bigr{\}}\\ ={}&\operatorname{span}\bigl{\{}\begin{bmatrix}r&Ap_{0}(A)r&\dotsi&Ap_{s-1}(A)r\end{bmatrix}\bigr{\}}\\ ={}&\operatorname{span}\{K_{1:s}\},\end{split}$ (50)
which also implies $\operatorname{span}\bigl{\{}\begin{bmatrix}r&AB_{1:s-1}\end{bmatrix}\bigr{\}}=\operatorname{span}\bigl{\{}\begin{bmatrix}r&AK_{1:s-1}\end{bmatrix}\bigr{\}}=\operatorname{span}\{V_{1:s}\}$.
Then assuming that these hold for $j-1$, i.e.,
$\operatorname{span}\{B_{1:(j-1)s}\}=\operatorname{span}\{K_{1:(j-1)s}\}=\operatorname{span}\{V_{1:(j-1)s}\},$ (51)
$\operatorname{span}\bigl{\{}\begin{bmatrix}r&AB_{1:(j-1)s-1}\end{bmatrix}\bigr{\}}=\operatorname{span}\bigl{\{}\begin{bmatrix}r&AK_{1:(j-1)s-1}\end{bmatrix}\bigr{\}}=\operatorname{span}\{V_{1:(j-1)s}\},$ (52)
we aim to prove that these hold for $j$. Recalling Algorithm 3, $B_{(j-1)s+1:js}$ is the $Q$-factor of $(I-V_{1:(j-1)s}V_{1:(j-1)s}^{\top})^{2}K_{(j-1)s+1:js}$ in exact arithmetic. Thus, we derive
$\begin{split}\operatorname{span}\{B_{1:js}\}={}&\operatorname{span}\{\begin{bmatrix}B_{1:(j-1)s}&B_{(j-1)s+1:js}\end{bmatrix}\}\\ ={}&\operatorname{span}\{\begin{bmatrix}V_{1:(j-1)s}&(I-V_{1:(j-1)s}V_{1:(j-1)s}^{\top})^{2}K_{(j-1)s+1:js}\end{bmatrix}\}\\ ={}&\operatorname{span}\{\begin{bmatrix}V_{1:(j-1)s}&K_{(j-1)s+1:js}\end{bmatrix}\}\\ ={}&\operatorname{span}\{K_{1:js}\}.\end{split}$ (53)
By $(I-V_{1:(j-1)s}V_{1:(j-1)s}^{\top})^{2}K_{(j-1)s+1:js-1}=B_{(j-1)s+1:js-1}S_{1:js-1,1:js-1}$, we similarly have $\operatorname{span}\{B_{1:js-1}\}=\operatorname{span}\{K_{1:js-1}\}$ and further from $\begin{bmatrix}r&AB_{1:js-1}\end{bmatrix}=V_{1:js}R_{1:js,1:js}$
$\operatorname{span}\{V_{1:js}\}=\operatorname{span}\bigl{\{}\begin{bmatrix}r&AB_{1:js-1}\end{bmatrix}\bigr{\}}=\operatorname{span}\bigl{\{}\begin{bmatrix}r&AK_{1:js-1}\end{bmatrix}\bigr{\}}.$ (54)
Thus, by (52) and (54), we obtain
$\begin{split}&\operatorname{span}\{V_{1:js}\}\\ &\quad=\operatorname{span}\bigl{\{}\begin{bmatrix}r&AK_{1:js-1}\end{bmatrix}\bigr{\}}\\ &\quad=\operatorname{span}\{V_{1:(j-1)s},AK_{(j-1)s:js-1}\}\\ &\quad=\operatorname{span}\{V_{1:(j-1)s},AK_{(j-1)s},Ap_{0}(A)V_{(j-1)s+1},\dotsc,Ap_{s-2}(A)V_{(j-1)s+1}\}\\ &\quad=\operatorname{span}\{V_{1:(j-1)s},V_{(j-1)s+1},Ap_{0}(A)V_{(j-1)s+1},\dotsc,Ap_{s-2}(A)V_{(j-1)s+1}\}\\ &\quad=\operatorname{span}\{K_{1:(j-1)s},K_{(j-1)s+1:js}\}\\ &\quad=\operatorname{span}\{K_{1:js}\}.\end{split}$ (55)
Thus, by induction on $j$, we draw the conclusion (49). ∎
As described in Theorem 1, $\kappa(\tilde{B}_{1:is})$, i.e., $\kappa(\hat{B}_{1:is}D_{1:is})$ for any diagonal $D_{1:is}$ with positive elements, has a critical influence on the backward error. Note that we perform an extra QR factorization aiming to make $B_{1:is}$ nearly orthonormal. Thus, we consider $D_{1:is}$ to be the identity matrix here. In the following lemma, we give some inspiration to show why $\kappa(\hat{B}_{1:is})$ can be expected to be well-conditioned regarding rounding errors. The proof of the lemma can be found in Appendix C.
Lemma 6.
Assuming that $V_{1:ks}$ is the exact result of Algorithm 2 with Algorithm 3, then there exists $Y_{(i-1)s+1:is}$ such that
$B_{(i-1)s+1:is}=V_{(i-1)s+1:is}Y_{(i-1)s+1:is},\quad\forall i\leq k.$ (56)
Furthermore, assume that $\hat{B}_{(i-1)s+1:is}$ and $\hat{V}_{(i-1)s+1:is}$ are the computed results of Algorithm 2 with Algorithm 3. If there exists a small perturbation $\Delta B_{(i-1)s+1:is}$ for any $i\leq k$ such that
$\hat{B}_{(i-1)s+1:is}+\Delta B_{(i-1)s+1:is}=\hat{V}_{(i-1)s+1:is}\tilde{Y}_{(i-1)s+1:is},$ (57)
and it holds that
$\sum_{i=1}^{k}\omega_{B_{i}}+5\sqrt{ks}\,\sum_{i=1}^{k}\bigl{\|}\Delta B_{(i-1)s+1:is}\bigr{\rVert}_{\mathsf{F}}+7s\,\omega_{k}\leq\frac{1}{2},$ (58)
then $\kappa(\hat{B}_{1:ks})\leq 2\sqrt{n}+\sqrt{s},$ where $\omega_{B_{i}}$ satisfies $\bigl{\|}\hat{B}_{(i-1)s+1:is}^{\top}\hat{B}_{(i-1)s+1:is}-I\bigr{\rVert}_{\mathsf{F}}\leq\omega_{B_{i}}$.
7 Numerical Experiments
In this section, we present numerical experiments to show that Algorithm 2 with Algorithm 3 ($s$-step GMRES with the modified $s$-step Arnoldi process) can employ a much larger block size $s$ compared to using the classical $s$-step Arnoldi process in Algorithm 1. All tests are performed in MATLAB R2023a.
7.1 Experiment settings
The following variants of the $s$-step GMRES algorithms are tested:
  1. 1.
    Classical $s$-step GMRES: $s$-step GMRES (Algorithm 2) with the classical $s$-step Arnoldi process (Algorithm 1), with stopping criterion (45).
  2. 2.
    Modified $s$-step GMRES with the additional criterion: $s$-step GMRES (Algorithm 2) with the modified Arnoldi process (Algorithm 3), with stopping criteria (45) and (46).
  3. 3.
    Classical $s$-step GMRES with the additional criterion: $s$-step GMRES (Algorithm 2) with the classical Arnoldi process (Algorithm 1), with stopping criteria (45) and (46).
In these three variants, BCGSI+ is employed as the orthogonalization method, respectively, in Line 4 of Algorithm 1 and in Line 5 of Algorithm 3. Since the condition number of the monomial basis grows exponentially with $s$, the Newton and Chebyshev bases are used to generate each sub-matrix defined in (48). To clearly demonstrate the theoretical results, we do not consider restarting and preconditioners in our numerical experiments, but we reiterate that one could extend the theoretical results to restarted $s$-step GMRES. The thresholds tol in (45) and tolH in (46) are set to, respectively, $n\bm{u}$ and $\sqrt{n}\bm{u}$.
For constructing the linear systems $Ax=b$, we choose three sparse square matrices, shown in Table 1, from the SuiteSparse Matrix Collection [^1] https://sparse.tamu.edu and the Matrix Market [^2] https://math.nist.gov/MatrixMarket as $A$. The two with larger condition numbers from the Matrix Market are suggested as test problems in [6]. The other matrix from the SuiteSparse Matrix Collection is a relatively well-conditioned matrix. The right-hand vector $b$ is set to be the vector of all ones and the initial guess $x_{0}=0$.
Table 1: Properties of test matrices: the condition number in this table is estimated by the MATLAB command svd.
Name Size Condition number
494_bus 0494 $2.42\times 10^{6}$0
fs1836 0183 $1.74\times 10^{11}$
sherman2 1,080 $9.64\times 10^{11}$
7.2 Tests for different block size $s$
We illustrate how varying the block size $s$ impacts the relative backward error $\frac{\bigl{\|}Ax-b\bigr{\rVert}}{\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\bigl{\|}x\bigr{\rVert}+\bigl{\|}b\bigr{\rVert}}$ and the iteration count, as computed by the different $s$-step GMRES variants, in Figures 2, 3, and 4.
From these figures, it is clear that the classical $s$-step GMRES algorithm must use very small values of $s$ to achieve satisfactory accuracy, since the condition number of $B_{1:p}$ increases rapidly with larger $s$. It is further clear that the modified $s$-step GMRES algorithm benefits from the well-conditioned basis $B_{1:p}$ and can utilize significantly larger $s$ in practice without sacrificing accuracy. It should be noted that until a point (depending on the matrix sparsity structure and machine parameters), a larger $s$ results in lower communication cost, but it usually necessitates more iterations to reach a given level of backward error. Therefore, it may not be beneficial in practice to set $s$ too large, even if it appears that a larger $s$ does not affect the backward error.
Figure 2: Relative backward errors (left) and the number of iterations (right) related to the block size $s$, computed by different $s$-step GMRES algorithms, for 494bus.
Figure 3: Relative backward errors (left) and the number of iterations (right) related to the block size $s$, computed by different $s$-step GMRES algorithms, for fs1836.
Figure 4: Relative backward errors (left) and the number of iterations (right) related to the block size $s$, computed by different $s$-step GMRES algorithms, for sherman2.
Then we choose three specific values of $s$, i.e., $s=1$, $4$, $16$, to show the behavior of the backward error related to the iteration in Figures 5, 6, and 7. As Lemma 5 predicts, there is almost no chance to obtain a better solution after the “key dimension” is reached; Figures 57 illustrate that using the criterion (46) achieves the ”optimal” accuracy for most cases. Thus, it is necessary to employ (46) to test the “key dimension”.
Figure 5: Relative backward errors by different $s$-step GMRES algorithms for 494bus with $s=1$, $4$, $16$, respectively, from left to right.
Figure 6: Relative backward errors by different $s$-step GMRES algorithms for fs1836 with $s=1$, $4$, $16$, respectively, from left to right.
Figure 7: Relative backward errors by different $s$-step GMRES algorithms for sherman2 with $s=1$, $4$, $16$, respectively, from left to right.
8 Conclusions
In this work, we provide an abstract framework for analyzing the backward stability of the preconditioned $s$-step GMRES algorithm. This framework accommodates various polynomial bases and block orthogonalization methods, demonstrating that the backward error of $s$-step GMRES is largely influenced by the condition number of the basis $B$. Moreover, it separately identifies the errors arising from the orthogonalization process (Step 2 in Section 3) and from solving the least squares problem (Step 3), facilitating its application to $s$-step GMRES with different block orthogonalization techniques. For standard GMRES, where $s=1$, our framework is an improved version of the modular GMRES framework introduced by [5]. We then apply the framework to analyze $s$-step GMRES with three widely-used block orthogonalization methods including block Householder QR, BMGS, and BCGSI+. Furthermore, based on the framework, we give the stopping criteria and discuss the requirements of the orthogonalization methods used in GMRES, illustrating why the MGS, CGSI+, Householder QR, and TSQR algorithms are often used for orthogonalization in GMRES.
We then provide an example, specifically Example 1, to illustrate the limitations of the classical $s$-step Arnoldi process (Algorithm 1). In certain scenarios, the condition number of the basis $B$ cannot be effectively managed by adaptively selecting $s$ to control the condition number of each sub-block of the basis, utilizing Newton or Chebyshev bases, or a restart process. Consequently, $s$-step GMRES may fail to achieve a satisfactory backward error even with a relatively small $s$. To address this issue, we introduce a modified $s$-step Arnoldi process that incorporates an extra QR factorization to make the basis $B$ well-conditioned. Numerical experiments demonstrate that this modified $s$-step Arnoldi process allows the use of a significantly larger $s$ while achieving the required accuracy. While the modified approach increases the communication cost per iteration by a factor of 2, it may be beneficial in cases where a higher $s$ can lead to greater per-iteration speedup. Future work involves high performance implementations on large-scale problems in order to better evaluate these tradeoffs.
Acknowledgments
Both authors are supported by the European Union (ERC, inEXASCALE, 101075632). Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. The first author additionally acknowledges support from the Charles University Research Centre program No. UNCE/24/SCI/005.
Appendix A Properties of BCGSI+
Given $X\in\mathbb{R}^{m\times n}$ with $m\geq n$, in Algorithm 4, we present the $(k+1)$-th step of the BCGSI+ algorithm to compute $X=QT$ with an orthonormal matrix $Q\in\mathbb{R}^{m\times n}$ and an upper triangular matrix $T\in\mathbb{R}^{n\times n}$. Note that MGS or any unconditionally stable QR algorithm, e.g., Householder QR or Tall-Skinny QR (TSQR), described in [10], can be utilized in Line 3, while any backward stable QR algorithm, i.e.,
$\hat{W}^{(2)}+\Delta W^{(2)}=\hat{Q}_{ks+1:(k+1)s}\hat{T}^{(2)}\quad\text{with}\quad\bigl{\|}\Delta W^{(2)}\bigr{\rVert}\leq O(\bm{u})\bigl{\|}\hat{W}^{(2)}\bigr{\rVert},$
can be employed in Line 6; see [7] for details.
0:  A matrix $X\in\mathbb{R}^{m\times n}$, the block size $s$, the orthogonal matrix $Q_{1:ks}\in\mathbb{R}^{m\times ks}$ satisfying $X_{1:ks}=Q_{1:ks}T_{1:ks,1:ks}$.
0:  The orthogonal matrix $Q_{1:(k+1)s}\in\mathbb{R}^{m\times(k+1)s}$ satisfies that $X_{1:(k+1)s}=Q_{1:(k+1)s}T_{1:(k+1)s}$.
1:  $S^{(1)}=Q_{1:ks}^{\top}X_{ks+1:(k+1)s}\in\mathbb{R}^{m\times s}$.
2:  $W^{(1)}=X_{ks+1:(k+1)s}-Q_{1:ks}S^{(1)}$.
3:  Compute $U\in\mathbb{R}^{m\times s}$ by QR algorithm such that $W^{(1)}=UT^{(1)}$.
4:  $S^{(2)}=Q_{1:ks}^{\top}U\in\mathbb{R}^{m\times s}$.
5:  $W^{(2)}=U-Q_{1:ks}S^{(2)}$.
6:  Compute $Q_{ks+1:(k+1)s}\in\mathbb{R}^{m\times s}$ by QR algorithm such that $W^{(2)}=Q_{ks+1:(k+1)s}T^{(2)}$.
7:  $T_{1:ks,ks+1:(k+1)s}=S^{(1)}+S^{(2)}T^{(1)}$.
8:  $T_{ks+1:(k+1)s,ks+1:(k+1)s}=T^{(2)}T^{(1)}$.
Algorithm 4 The $(k+1)$-th step of the BCGSI+ algorithm
The properties of BCGSI+ have already been studied in [7, Section 2.3]. Based on the results in [7, Section 2.3], it is easy to obtain the following lemmas.
Lemma 7.
Let $\widehat{Q}_{1:js}$ and $\widehat{T}_{1:js}$ be computed by Algorithm 4. If assuming $O(\bm{u})\kappa(X_{1:js})<1$, then
$X_{1:js}+\Delta X_{1:js}=\widehat{Q}_{1:js}\widehat{T}_{1:js},\quad\bigl{\|}\Delta X_{i}\bigr{\rVert}\leq O(\bm{u})\bigl{\|}X_{i}\bigr{\rVert}$ (59)
for any $i\leq js$, and
$\bigl{\|}\widehat{Q}_{1:js}^{\top}\widehat{Q}_{1:js}-I\bigr{\rVert}_{\mathsf{F}}\leq O(\bm{u}).$ (60)
Proof.
Similarly to the proof of [7, Theorem 2], it is easy to verify (59), since each line of Algorithm 4 is column-wise backward stable. The conclusion (60) is directly followed by [7, Theorem 2]. ∎
From Lemma 7, $\omega_{\mathrm{qr}}=O(\bm{u})$, and (13) is satisfied when the $Q$-factor is not well-conditioned. Then it remains to estimate $\omega_{H}$ defined by (26).
Lemma 8.
Assume that for $is+j$, $\widehat{Q}_{1:is+j}$ and $\widehat{T}_{1:is+j}$ are computed by Algorithm 4. If
$\bigl{\|}\widehat{Q}_{1:is+j-1}^{\top}\widehat{Q}_{1:is+j-1}-I\bigr{\rVert}_{\mathsf{F}}\leq O(\bm{u}),$ (61)
$\bigl{\|}\widehat{Q}_{1:is+j}^{\top}\widehat{Q}_{1:is+j}-I\bigr{\rVert}_{\mathsf{F}}>O(\bm{u}),$ (62)
then
$\max\{\bigl{\|}\widehat{Q}_{is+j}\widehat{T}_{is+j,is+j}\bigr{\rVert},\bigl{\lvert}\widehat{T}_{is+j,is+j}\bigr{\rvert}\}\leq O(\bm{u})\bigl{\|}X_{is+j}\bigr{\rVert}.$ (63)
Proof.
Without loss of generality, we only need to prove the case $j=1$ since Algorithm 4 is columnwise backward stable. From the assumption (61),
$\bigl{\|}\widehat{Q}_{1:is+1}^{\top}\widehat{Q}_{1:is+1}-I\bigr{\rVert}_{\mathsf{F}}\leq O(\bm{u})+\bigl{\|}\widehat{Q}_{is+1}^{\top}\widehat{Q}_{is+1}-I\bigr{\rVert}_{\mathsf{F}}+2\bigl{\|}\widehat{Q}_{1:is}^{\top}\widehat{Q}_{is+1}\bigr{\rVert}_{\mathsf{F}}.$ (64)
Note that [7, Equations (41)–(43), (46), (47), and (56)] do not depend on [7, Assumption (40)] when using an unconditionally stable QR algorithm in Line 3 of Algorithm 4. From [7, Equation (47)] it holds that
$\bigl{\|}\widehat{Q}_{is+1}^{\top}\widehat{Q}_{is+1}-I\bigr{\rVert}_{\mathsf{F}}\leq O(\bm{u}).$
Then it remains to estimate $\bigl{\|}\widehat{Q}_{1:is}^{\top}\widehat{Q}_{is+1}\bigr{\rVert}_{\mathsf{F}}$, which can be bounded as
$\begin{split}\bigl{\|}\widehat{Q}_{1:is}^{\top}\widehat{Q}_{is+1}\bigr{\rVert}_{\mathsf{F}}&\leq\bigl{\|}\widehat{Q}_{1:is}^{\top}\tilde{W}^{(2)}_{1}(\widehat{T}^{(2)}_{1,1})^{-1}\bigr{\rVert}_{\mathsf{F}}+\bigl{\|}\widehat{Q}_{1:is}^{\top}\Delta\tilde{W}^{(2)}_{1}(\widehat{T}^{(2)}_{1,1})^{-1}\bigr{\rVert}_{\mathsf{F}}\\ &\leq\frac{\bigl{\|}(I-\widehat{Q}_{1:is}^{\top}\widehat{Q}_{1:is})\widehat{Q}_{1:is}^{\top}\widehat{U}_{1}\bigr{\rVert}_{\mathsf{F}}}{\bigl{\lvert}\widehat{T}^{(2)}_{1,1}\bigr{\rvert}}+\frac{\bigl{\|}\widehat{Q}_{1:is}\bigr{\rVert}_{\mathsf{F}}\bigl{\|}\Delta\tilde{W}^{(2)}_{1}\bigr{\rVert}}{\bigl{\lvert}\widehat{T}^{(2)}_{1,1}\bigr{\rvert}},\end{split}$ (65)
where $\widehat{W}^{(2)}=\tilde{W}^{(2)}+\Delta\tilde{W}^{(2)}$ with $\tilde{W}^{(2)}=(I-\widehat{Q}_{1:is}\widehat{Q}_{1:is}^{\top})\widehat{U}$. Together with [7, Equations (43) and (56)] and the assumption (61), we obtain
$\bigl{\|}\widehat{Q}_{1:is}^{\top}\widehat{Q}_{is+1}\bigr{\rVert}_{\mathsf{F}}\leq\frac{O(\bm{u})}{\bigl{\lvert}\widehat{T}^{(2)}_{1,1}\bigr{\rvert}}.$ (66)
Then together with [7, Lemmas 2 and 6], it follows that
$\tilde{W}^{(2)}_{1}+\Delta W^{(2)}_{1}=\widehat{Q}_{is+1}\widehat{T}^{(2)}_{1,1},\quad\bigl{\|}\Delta W^{(2)}_{1}\bigr{\rVert}\leq O(\bm{u}),$
and further,
$\begin{split}\bigl{\|}\widehat{Q}_{1:is}^{\top}\widehat{Q}_{is+1}\bigr{\rVert}_{\mathsf{F}}&\leq\frac{O(\bm{u})\bigl{\|}\widehat{Q}_{is+1}\bigr{\rVert}}{\bigl{\|}(I-\widehat{Q}_{1:is}\widehat{Q}_{1:is}^{\top})\widehat{U}_{1}\bigr{\rVert}-\bigl{\|}\Delta W^{(2)}_{1}\bigr{\rVert}}\\ &\leq\frac{O(\bm{u})}{\bigl{\|}(I-\widehat{Q}_{1:is}\widehat{Q}_{1:is}^{\top})\widehat{U}_{1}\bigr{\rVert}-O(\bm{u})}.\end{split}$
This means that $\bigl{\|}\widehat{Q}_{1:is+1}^{\top}\widehat{Q}_{1:is+1}-I\bigr{\rVert}_{\mathsf{F}}\leq O(\bm{u})$ if $2\bigl{\|}(I-\widehat{Q}_{1:is}\widehat{Q}_{1:is}^{\top})\widehat{U}_{1}\bigr{\rVert}>O(\bm{u})$ holds. Furthermore, the contrapositive is that $2\bigl{\|}(I-\widehat{Q}_{1:is}\widehat{Q}_{1:is}^{\top})\widehat{U}_{1}\bigr{\rVert}\leq O(\bm{u})$ if $\bigl{\|}\widehat{Q}_{1:is+1}^{\top}\widehat{Q}_{1:is+1}-I\bigr{\rVert}_{\mathsf{F}}>O(\bm{u})$ guaranteed by the assumption (62). Then by [7, Lemma 6], and
$\widehat{T}_{is+1,is+1}=\widehat{T}^{(2)}_{1,1}\widehat{T}^{(1)}_{1,1}+\Delta T_{is+1,is+1}\quad\text{with}\quad\bigl{\lvert}\Delta T_{is+1,is+1}\bigr{\rvert}\leq O(\bm{u})\bigl{\lvert}\widehat{T}^{(2)}_{1,1}\bigr{\rvert}\bigl{\lvert}\widehat{T}^{(1)}_{1,1}\bigr{\rvert},$
we have
$\begin{split}\bigl{\|}\widehat{Q}_{is+1}\widehat{T}_{is+1,is+1}\bigr{\rVert}\leq\bigl{\|}\widehat{Q}_{is+1}\widehat{T}^{(2)}_{1,1}\bigr{\rVert}\bigl{\lvert}\widehat{T}^{(1)}_{1,1}\bigr{\rvert}+O(\bm{u})\bigl{\|}\widehat{Q}_{is+1}\bigr{\rVert}\bigl{\lvert}\widehat{T}^{(2)}_{1,1}\bigr{\rvert}\bigl{\lvert}\widehat{T}^{(1)}_{1,1}\bigr{\rvert}\leq O(\bm{u})\bigl{\|}X_{is+1}\bigr{\rVert},\end{split}$
which also implies $\bigl{\lvert}\widehat{T}_{is+1,is+1}\bigr{\rvert}\leq O(\bm{u})\bigl{\|}X_{is+1}\bigr{\rVert}$ by noticing $\bigl{\|}\widehat{Q}_{is+1}\bigr{\rVert}\geq 1-O(\bm{u})$. ∎
Appendix B Example for classical $s$-step GMRES with BCGSPIPI+
Example 2.
We construct the linear system $Ax=b$, where $A$ is a $20$-by-$20$ random matrix with $\kappa(A)=10^{10}$ generated using the MATLAB command rng(1) and gallery(’randsvd’, [20, 20], 1e10, 5). The vector $b$ is selected as the right singular vector corresponding to the fourth largest singular value, and the initial guess $x_{0}$ is the zero vector.
For this specific linear system, the relative backward error $\frac{\bigl{\|}Ax-b\bigr{\rVert}}{\bigl{\|}A\bigr{\rVert}_{\mathsf{F}}\bigl{\|}x\bigr{\rVert}+\bigl{\|}b\bigr{\rVert}}$ of the solution computed by using standard GMRES with CGSPIPI+, namely BCGSPIPI+ with $s=1$, is approximately $10^{-8}$. In contrast, using standard GMRES with CGSI+ results in an error of around $10^{-16}$, as illustrated in Figure 8. This difference occurs because CGSPIPI+ cannot generate a nearly orthonormal basis $\hat{V}_{1:20}$ when the condition number of $\begin{bmatrix}\hat{r}&\hat{W}_{1:19}\end{bmatrix}$ exceeds approximately $10^{8}$, implying $O(\bm{u})\kappa^{2}(\begin{bmatrix}\hat{r}&\hat{W}_{1:19}\end{bmatrix})>1$. Under this situation, the return value of $\hat{V}_{20}$ is a NaN vector. Thus, it is not possible to obtain a more accurate solution than $x^{(19)}$, whose backward error is approximately $10^{-8}$.
A similar result occurs for $s$-step GMRES with $s=2$. The relative backward error using BCGSPIPI+ is approximately $10^{-5}$, while for $s$-step GMRES with BCGSI+, the error is around $10^{-11}$, as illustrated in Figure 9.
Figure 8: The plot for Example 2: The plot is of relative backward error. Each line with different color in the plot denotes the standard GMRES using different orthogonalization methods, including CGSI+ and CGSPIPI+.
Figure 9: The plot for Example 2: From left to right, the plots are of relative backward error and the condition number of the basis $\tilde{B}$, where we normalize each column of $\hat{B}$ as $\tilde{B}$. Each line with different color in the plots denotes the classical $s$-step GMRES with BCGSI+ or BCGSPIPI+ using different polynomials, including Newton and Chebyshev polynomials, to generate the basis by (48).
Appendix C Proof of Lemma 6
Proof of Lemma 6.
First, we aim to prove (56) by induction. For the base case, from (49) with $i=1$, there exists $Y_{1:s}$ such that $B_{1:s}=V_{1:s}Y_{1:s}$. Assume that $B_{(i-1)s+1:is}=V_{(i-1)s+1:is}Y_{(i-1)s+1:is}$ holds for all $i\leq j-1$. Then our aim is to prove that it holds for $j$. Since $V_{1:(j-1)s}$ is orthonormal and $B_{(j-1)s+1:js}$ is the $Q$-factor of $(I-V_{1:(j-1)s}V_{1:(j-1)s}^{\top})^{2}K_{(j-1)s+1:js}$, we obtain $V_{1:(j-1)s}^{\top}B_{(j-1)s+1:js}=0$. Together with the above assumptions on $i\leq j-1$ and (49), there exists $Y_{(j-1)s+1:js}$ such that
$B_{1:js}=V_{1:js}\begin{bmatrix}Y_{1:(j-1)s}&V_{1:(j-1)s}^{\top}B_{(j-1)s+1:js}\\ 0&Y_{(j-1)s+1:js}\end{bmatrix}=V_{1:js}\begin{bmatrix}Y_{1:(j-1)s}&0\\ 0&Y_{(j-1)s+1:js}\end{bmatrix},$
which gives (56) by induction on $j$.
Then we will bound $\kappa(\hat{B}_{1:ks})$. By the definition of $\omega_{B_{i}}$, we only need to consider the off-diagonal blocks $\bigl{\|}\hat{B}_{(i-1)s+1:is}^{\top}\hat{B}_{(j-1)s+1:js}\bigr{\rVert}_{\mathsf{F}}$. From (57) and dropping the quadratic terms, it holds that
$\begin{split}&\bigl{\|}\hat{B}_{(i-1)s+1:is}^{\top}\hat{B}_{(j-1)s+1:js}\bigr{\rVert}_{\mathsf{F}}\\ &\quad\leq\bigl{\|}\Delta B_{(i-1)s+1:is}\bigr{\rVert}_{\mathsf{F}}\bigl{\|}\hat{V}_{(j-1)s+1:js}\bigr{\rVert}\bigl{\|}\tilde{Y}_{(j-1)s+1:js}\bigr{\rVert}\\ &\qquad+\bigl{\|}\tilde{Y}_{(i-1)s+1:is}\bigr{\rVert}\bigl{\|}\hat{V}_{(i-1)s+1:is}^{\top}\hat{V}_{(j-1)s+1:js}\bigr{\rVert}_{\mathsf{F}}\bigl{\|}\tilde{Y}_{(j-1)s+1:js}\bigr{\rVert}\\ &\qquad+\bigl{\|}\Delta B_{(j-1)s+1:js}\bigr{\rVert}_{\mathsf{F}}\bigl{\|}\hat{V}_{(i-1)s+1:is}\bigr{\rVert}\bigl{\|}\tilde{Y}_{(i-1)s+1:is}\bigr{\rVert}\\ &\quad\leq\bigl{\|}\Delta B_{(i-1)s+1:is}\bigr{\rVert}_{\mathsf{F}}\bigl{\|}\tilde{Y}_{(j-1)s+1:js}\bigr{\rVert}\\ &+\bigl{\|}\tilde{Y}_{(i-1)s+1:is}\bigr{\rVert}\bigl{\|}\hat{V}_{(i-1)s+1:is}^{\top}\hat{V}_{(j-1)s+1:js}\bigr{\rVert}_{\mathsf{F}}\bigl{\|}\tilde{Y}_{(j-1)s+1:js}\bigr{\rVert}\\ &\qquad+\bigl{\|}\Delta B_{(j-1)s+1:js}\bigr{\rVert}_{\mathsf{F}}\bigl{\|}\tilde{Y}_{(i-1)s+1:is}\bigr{\rVert},\end{split}$
which implies that
$\begin{split}\sum_{i,j=1;i\neq j}^{k}&\bigl{\|}\hat{B}_{(i-1)s+1:is}^{\top}\hat{B}_{(j-1)s+1:js}\bigr{\rVert}_{\mathsf{F}}^{2}\\ \leq{}&3\sum_{i,j=1;i\neq j}^{k}\bigl{(}\bigl{\|}\Delta B_{(i-1)s+1:is}\bigr{\rVert}_{\mathsf{F}}^{2}\bigl{\|}\tilde{Y}_{(j-1)s+1:js}\bigr{\rVert}^{2}\\ &+\bigl{\|}\tilde{Y}_{(i-1)s+1:is}\bigr{\rVert}^{2}\bigl{\|}\hat{V}_{(i-1)s+1:is}^{\top}\hat{V}_{(j-1)s+1:js}\bigr{\rVert}_{\mathsf{F}}^{2}\bigl{\|}\tilde{Y}_{(j-1)s+1:js}\bigr{\rVert}^{2}\\ &+\bigl{\|}\Delta B_{(j-1)s+1:js}\bigr{\rVert}_{\mathsf{F}}^{2}\bigl{\|}\tilde{Y}_{(i-1)s+1:is}\bigr{\rVert}^{2}\bigr{)}\\ \leq{}&6k\sum_{i=1}^{k}\bigl{(}\bigl{\|}\Delta B_{(i-1)s+1:is}\bigr{\rVert}_{\mathsf{F}}^{2}\bigl{\|}\tilde{Y}_{(j-1)s+1:js}\bigr{\rVert}^{2}\bigr{)}+3\max_{i}\bigl{(}\bigl{\|}\tilde{Y}_{(j-1)s+1:js}\bigr{\rVert}^{4}\bigr{)}\omega_{k}^{2}.\end{split}$
Together with $\bigl{\|}\hat{B}_{(i-1)s+1:is}\bigr{\rVert}_{\mathsf{F}}\leq\sqrt{s}(1+\omega_{B_{i}})$ and
$\bigl{\|}\tilde{Y}_{(i-1)s+1:is}\bigr{\rVert}\leq\frac{\bigl{\|}\hat{B}_{(i-1)s+1:is}\bigr{\rVert}_{\mathsf{F}}+\bigl{\|}\Delta B_{(i-1)s+1:is}\bigr{\rVert}_{\mathsf{F}}}{1-\bigl{\|}\hat{V}_{(i-1)s+1:is}^{\top}\hat{V}_{(i-1)s+1:is}-I\bigr{\rVert}_{\mathsf{F}}},$ (67)
we obtain
$\begin{split}&\bigl{\|}\hat{B}_{1:ks}^{\top}\hat{B}_{1:ks}-I\bigr{\rVert}_{\mathsf{F}}^{2}\\ &\quad=\sum_{i=1}^{k}\bigl{\|}\hat{B}_{(i-1)s+1:is}^{\top}\hat{B}_{(i-1)s+1:is}-I\bigr{\rVert}_{\mathsf{F}}^{2}+\sum_{i,j=1;i\neq j}^{k}\bigl{\|}\hat{B}_{(i-1)s+1:is}^{\top}\hat{B}_{(j-1)s+1:js}\bigr{\rVert}_{\mathsf{F}}^{2}\\ &\quad\leq\sum_{i=1}^{k}\omega_{B_{i}}^{2}+6k\sum_{i=1}^{k}\bigl{(}\bigl{\|}\Delta B_{(i-1)s+1:is}\bigr{\rVert}_{\mathsf{F}}^{2}\bigl{\|}\tilde{Y}_{(j-1)s+1:js}\bigr{\rVert}^{2}\bigr{)}+3\max_{i}\bigl{(}\bigl{\|}\tilde{Y}_{(j-1)s+1:js}\bigr{\rVert}^{4}\bigr{)}\omega_{k}\\ &\quad\leq\sum_{i=1}^{k}\omega_{B_{i}}^{2}+24ks\sum_{i=1}^{k}\bigl{(}\bigl{\|}\Delta B_{(i-1)s+1:is}\bigr{\rVert}_{\mathsf{F}}^{2}\bigr{)}+48s^{2}\omega_{k}^{2}\end{split}$
by dropping the quadratic terms. This implies that, from the assumption (58),
$\begin{split}\bigl{\|}\hat{B}_{1:ks}^{\top}\hat{B}_{1:ks}-I\bigr{\rVert}_{\mathsf{F}}\leq\sum_{i=1}^{k}\omega_{B_{i}}+5\sqrt{ks}\sum_{i=1}^{k}\bigl{\|}\Delta B_{(i-1)s+1:is}\bigr{\rVert}_{\mathsf{F}}+7s\,\omega_{k}\leq\frac{1}{2}.\end{split}$ (68)
Thus, $\sigma_{\min}(\hat{B}_{1:ks})$ can be bounded by
$\sigma_{\min}(\hat{B}_{1:ks})\geq 1-\bigl{\|}\hat{B}_{1:ks}^{\top}\hat{B}_{1:ks}-I\bigr{\rVert}_{\mathsf{F}}\geq\frac{1}{2}$ (69)
and further
$\kappa(\hat{B}_{1:ks})\leq\frac{\sum_{i}\bigl{\|}\hat{B}_{(i-1)s+1:is}\bigr{\rVert}_{\mathsf{F}}}{\sigma_{\min}(\hat{B}_{1:ks})}\leq 2\sqrt{\sum_{i=1}^{k}s\,(1+\omega_{B_{i}})}\leq 2\sqrt{n}+\sqrt{s}.$ (70)